STK Reference manual

Version 3.99

Erick Gallesio

Université de Nice - Sophia Antipolis Laboratoire I3S - CNRS URA 1376 - ESSI.

Route des Colles

B.P. 145

06903 Sophia-Antipolis Cedex - FRANCE

email: eg@unice.fr

April 1998 Document Reference

Erick Gallesio, STk Reference Manual, RT 95-31d, I3S-CNRS / Université de Nice - Sophia Antipolis, juillet 1995.

Contents

I Reference Manual 5

1 Overview of STk . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 7 2 Lexical conventions . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 7

2.1 Identifiers . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 7 2.2 Comments . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 7 2.3 Other notations . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 8

3 Basic concepts . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 8 4 Expressions . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 8

4.1 Primitive expression types . . . . . . . . . . . . . . . . . . . . . . . . . 8 4.2 Derived expression types . . . . . . . . . . . . . . . . . . . . . . . . . . 9

5 Program structure . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 11 6 Standard procedures . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 11

6.1 Booleans . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 11 6.2 Equivalence predicates . . . . . . . . . . . . . . . . . . . . . . . . . . . 11 6.3 Pairs and lists . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 12 6.4 Symbols . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 13 6.5 Numbers . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 14 6.6 Characters . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 17 6.7 Strings . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 18 6.8 Vectors . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 19 6.9 Control features . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 20 6.10 Input and output . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 21 6.11 Keywords . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 28 6.12 Tk commands . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 29 6.13 Modules . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 30 6.14 Environments . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 34 6.15 Macros . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 35 6.16 System procedures . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 36 6.17 Addresses . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 39 6.18 Signals . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 39 6.19 Hash tables . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 41 6.20 Regular expressions . . . . . . . . . . . . . . . . . . . . . . . . . . . . 44 6.21 Pattern matching . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 46 6.22 Processes . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 49 6.23 Sockets . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 51 6.24 Foreign Function Interface . . . . . . . . . . . . . . . . . . . . . . . . . 54 6.25 Miscellaneous . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 58

II Annexes 65

A Using the Tk toolkit 67 1 Calling a Tk-command . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 67 2 Associating Callbacks to Tk-commands . . . . . . . . . . . . . . . . . . . . . 68 3 Tk bindings . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 69

B Differences with R4RS 71 1 Symbols . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 71 2 Types . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 73 3 Procedures . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 73

C An introduction to STklos 75 1 Introduction . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 75 2 Class definition and instantiation . . . . . . . . . . . . . . . . . . . . . . . . . 75

2.1 Class definition . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 75 3 Inheritance . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 76

3.1 Class hierarchy and inheritance of slots . . . . . . . . . . . . . . . . . 76 3.2 Instance creation and slot access . . . . . . . . . . . . . . . . . . . . . 77 3.3 Slot description . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 78 3.4 Class precedence list . . . . . . . . . . . . . . . . . . . . . . . . . . . . 81

4 Generic functions . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 82 4.1 Generic functions and methods . . . . . . . . . . . . . . . . . . . . . . 82 4.2 Next-method . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 83 4.3 Example . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 84

D Modules: Examples 87

E Changes 91

F Miscellaneous Informations 95 1 Introduction . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 95 2 About STk . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 95

2.1 Latest release . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 95 2.2 Sharing Code . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 95 2.3 STk Mailing list . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 95 2.4 STk FAQ . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 96 2.5 Reporting a bug . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 96

3 STk and Emacs . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 96

3.1 Using the SLIB package with STk . . . . . . . . . . . . . . . . . . . . 97 4 Getting information about Scheme . . . . . . . . . . . . . . . . . . . . . . . . 97

4.1 The R^{4}RS document . . . . . . . . . . . . . . . . . . . . . . . . . . . . 97 4.2 The Scheme Repository . . . . . . . . . . . . . . . . . . . . . . . . . . 98 4.3 Usenet newsgroup and other addresses . . . . . . . . . . . . . . . . . . 98

Part I

Reference Manual

Introduction

This document provides a complete list of procedures and special forms implemented in version 3.99 of STk. Since STk is (nearly) compliant with the language described in the Revised^{4} Report on the Algorithmic Language Scheme (denoted R^{4}RS hereafter^{1})[1], the organization of this manual follows the R^{4}RS and only describes extensions.

1 Overview of STk

Today’s graphical toolkits for applicative languages are often not satisfactory. Most of the time, they ask the user to be an X window system expert and force him/her to cope with arcane details such as server connections and event queues. This is a real problem, since programmers using this kind of languages are generally not inclined to system programming, and few of them will bridge the gap between the different abstraction levels.

Tk is a powerful graphical toolkit promising to fill that gap. It was developed at the University of Berkeley by John Ousterhout [2]. The toolkit offers high level widgets such as buttons or menus and is easily programmable, requiring little knowledge of X fundamentals. Tk relies on an interpretative shell-like language named Tcl [3].

STk is an implementation of the Scheme programming language, providing a full integration of the Tk toolkit. In this implementation, Scheme establishes the link between the user and the Tk toolkit, replacing Tcl.

2 Lexical conventions

2.1 Identifiers

Syntactic keywords can be used as variables in STk. Users must be aware that this extension of the language could lead to ambiguities in some situations.

2.2 Comments

There are three types of comments in STk:

1. a semicolon (;) indicates the start of a comment. This kind of comment extends to the end of the line (as described in R^{4}RS).

2. multi-lines comment use the classical Lisp convention: a comment begins with #| and ends with |#.

3. comments can also be introduced by #!. This extension is particularly useful for building STk scripts. On most Unix implementations, if the first line of a script looks like this:

             #!/usr/local/bin/stk -file

1The Revised^{4} Report on the Algorithmic Language Scheme is available through anonymous FTP from ftp.cs.indiana.edu in the directory /pub/scheme-repository/doc

then the script can be started directly as if it were a binary. STk is loaded behind the scenes and reads and executes the script as a Scheme program. Of course this assumes that STk is located in /usr/local/bin.

2.3 Other notations

STk accepts all the notations defined in R^{4}RS plus

[ ] Brackets are equivalent to parentheses. They are used for grouping and to notate lists. A list opened with a left square bracket must be closed with a right square bracket (section 6.3).

: A colon at the beginning of a symbol introduces a keyword. Keywords are described in section 6.11.

#.<expr> is read as the evaluation of the Scheme expression <expr>. The evaluation is done during the read process, when the #. is encountered. Evaluation is done in the environment of the current module.

             (define foo 1)
             #.foo
                   ⟹ 1
             '(foo #.foo #.(+ foo foo))
                   ⟹ (foo 1 2)
             (let ((foo 2))
                #.foo)
                   ⟹ 1

3 Basic concepts

Identical to R^{4}RS.

4 Expressions

4.1 Primitive expression types

(quote ⟨datum⟩)  syntax
'⟨datum⟩  syntax

The quoting mechanism is identical to R^{4}RS. Keywords (see section 6.11), as numerical constants, string constants, character constants, and boolean constants evaluate “to themselves”; they need not be quoted.

    '"abc"  ⟹ "abc"
    "abc"  ⟹ "abc"
    '145932  ⟹ 145932
    145932  ⟹ 145932
    '#t  ⟹ #t
    #t  ⟹ #t

    ':key  ⟹ :key
    :key  ⟹ :key

Note: R^{4}RS requires to quote constant lists and constant vectors. This is not necessary with STk.

(⟨operator⟩ ⟨operand_{1}⟩ …)  syntax

Identical to R^{4}RS. Furthermore, ⟨operator⟩ can be a macro (see section 6.15).

(lambda ⟨formals⟩ ⟨body⟩)  syntax
(if ⟨test⟩ ⟨consequent⟩ ⟨alternate⟩)  syntax
(if ⟨test⟩ ⟨consequent⟩)  syntax
(set! ⟨variable⟩ ⟨expression⟩)  syntax

Identical to R^{4}RS.

4.2 Derived expression types

(cond ⟨clause_{1}⟩ ⟨clause_{2}⟩ …)  syntax
(case ⟨key⟩ ⟨clause_{1}⟩ ⟨clause_{2}⟩ …)  syntax
(and ⟨test_{1}⟩ …)  syntax
(or ⟨test_{1}⟩ …)  syntax

Identical to R^{4}RS.

(when ⟨test⟩ ⟨expression_{1}⟩ ⟨expression_{2}⟩ …)  syntax

If the ⟨test⟩ expression yields a true value, the ⟨expression⟩s are evaluated from left to right and the value of the last ⟨expression⟩ is returned.

(unless ⟨test⟩ ⟨expression_{1}⟩ ⟨expression_{2}⟩ …)  syntax

If the ⟨test⟩ expression yields a false value, the ⟨expression⟩s are evaluated from left to right and the value of the last ⟨expression⟩ is returned.

(let ⟨bindings⟩ ⟨body⟩)  syntax
(let ⟨variable⟩ ⟨bindings⟩ ⟨body⟩)  syntax
(let* ⟨bindings⟩ ⟨body⟩)  syntax

Identical to R^{4}RS.

(fluid-let ⟨bindings⟩ ⟨body⟩)  syntax

The bindings are evaluated in the current environment, in some unspecified order, the current values of the variables present in bindings are saved, and the new evaluated values are assigned to the bindings variables. Once this is done, the expressions of body are evaluated sequentially in the current environment; the value of the last expression is the result of fluid-let. Upon exit, the stored variables values are restored. An error is signalled if any of the bindings variable is unbound.

    (let* ((a 'out)
           (f (lambda () a)))
      (list a
            (fluid-let ((a 'in)) (f))
            a))
              ⟹ (out in out)

When the body of a fluid-let is exited by invoking a continuation, the new variable values are saved, and the variables are set to their old values. Then, if the body is reentered by invoking a continuation, the old values are saved and new values are restored. The following example illustrates this behaviour

    (let ((cont #f)
          (l    '())
          (a    'out))
      (set! l (cons a l))
      (fluid-let ((a 'in))
        (set! cont (call/cc (lambda (k) k)))
        (set! l (cons a l)))
      (set! l (cons a l))
      (if cont (cont #f) l))
              ⟹ (out in out in out)

(letrec ⟨bindings⟩ ⟨body⟩)  syntax
(begin ⟨expression_{1}⟩ ⟨expression_{2}⟩ …)  syntax
(do ⟨inits⟩ ⟨test⟩ ⟨body⟩)  syntax
(delay ⟨expression⟩)  syntax
(quasiquote ⟨template⟩)  syntax
` ⟨template⟩  syntax

Identical to R^{4}RS.

(dotimes (var count) ⟨expression_{1}⟩ ⟨expression_{2}⟩ …)  syntax
(dotimes (var count result) ⟨expression_{1}⟩ ⟨expression_{2}⟩ …)  syntax

Dotimes evaluates the count form, which must return an integer. It then evaluates the ⟨expression⟩s once for each integer from zero (inclusive) to count (exclusive), in order, with the variable var bound to the integer; if the value of count is zero or negative, then the ⟨expression⟩s are not evaluated. When the loop completes, result is evaluated and its value is returned as the value of the dotimes expression. If result is omitted, dotimes returns #f.

    (let ((l '()))
      (dotimes (i 4 l)
         (set! l (cons i l))))
              ⟹ (3 2 1 0)

(while ⟨test⟩ ⟨expression_{1}⟩ ⟨expression_{2}⟩ …)  syntax

While evaluates the ⟨expression⟩s until ⟨test⟩ returns a false value. The value of a while construct is unspecified.

(until ⟨test⟩ ⟨expression_{1}⟩ ⟨expression_{2}⟩ …)  syntax

Until evaluates the ⟨expression⟩s while ⟨test⟩ returns a false value. The value of an unless construct is unspecified.

5 Program structure

Identical to R^{4}RS.

6 Standard procedures

6.1 Booleans

In STk the boolean value #f is different from the empty list, as required by R^{4}RS.

(not obj )  procedure
(boolean? obj )  procedure

Identical to R^{4}RS.

6.2 Equivalence predicates

(eqv? obj_{1} obj_{2})  procedure

STk extends the eqv? predicate defined in the R^{4}RS to take keywords into account: if obj_{1} and obj_{2} are both keywords, the eqv? predicate will yield #t if and only if

    (string=? (keyword->string obj1)
              (keyword->string obj2))
                                ⟹ #t

(eq? obj_{1} obj_{2})  procedure

STk extends the eq? predicate defined in R^{4}RS to take keywords into account. On keywords, eq? behaves like eqv?.

    (eq? :key :key)  ⟹ #t

(equal? obj_{1} obj_{2})  procedure

Identical to R^{4}RS.

6.3 Pairs and lists

(pair? obj )  procedure
(cons obj_{1} obj_{2})  procedure
(car pair)  procedure
(cdr pair)  procedure
(set-car! pair obj )  procedure
(set-cdr! pair obj )  procedure
(caar pair)  procedure
(cadr pair)  procedure
⋮                                                                   ⋮
(cdddar pair)  procedure
(cddddr pair)  procedure
(null? obj )  procedure
(list? obj )  procedure
(list obj …)  procedure
(length list)  procedure
(append list …)  procedure

Identical to R^{4}RS.

(append! list …)  procedure

Returns a list consisting of the elements of the first list followed by the elements of the other lists, as with append. The differenece with append is that the arguments are changed rather than copied.

      (append! '(1 2) '(3 4) '(5 6))
              ⟹ '(1 2 3 4 5 6)
      (let ((l1 '(1 2))
            (l2 '(3 4))
            (l3 '(5 6)))
        (append! l1 l2 l3)
        (list l1 l2 l3))
              ⟹ ((1 2 3 4 5 6) (3 4 5 6) (5 6))

(reverse list)  procedure
(list-tail list k)  procedure
(list-ref list k)  procedure
(memq obj list)  procedure
(memv obj list)  procedure
(member obj list)  procedure
(assq obj alist)  procedure
(assv obj alist)  procedure
(assoc obj alist)  procedure

Identical to R^{4}RS.

(remq obj list)  procedure
(remv obj list)  procedure
(remove obj list)  procedure

Each function return a copy of list where all the occurences of obj have been deleted. The predicate used to test the presence of obj in list is respectively eq, eqv and equal.

Note: It is not an error if obj does not appear in list.

    (remq 1 '(1 2 3))  ⟹ (2 3)
    (remq "foo" '("foo" "bar")) ⟹ ("foo" "bar")
    (remove "foo" '("foo" "bar"))
                                ⟹ ("bar")

(last-pair list)  procedure

Returns the last pair of list^{2}.

    (last-pair '(1 2 3))
              ⟹ 3
    (last-pair '(1 2 . 3))
              ⟹ (2 . 3)

(list* obj )  procedure

list* is like list except that the last argument to list* is used as the cdr of the last pair constructed.

    (list* 1 2 3)  ⟹ (1 2 . 3)
    (list* 1 2 3 '(4 5))  ⟹ (1 2 3 4 5)

(copy-tree obj )  procedure

Copy-tree recursively copies trees of pairs. If obj is not a pair, it is returned; otherwise the result is a new pair whose car and cdr are obtained by calling copy-tree on the car and cdr of obj, respectively.

6.4 Symbols

The STk reader can cope with symbols whose names contain special characters or letters in the non standard case. When a symbol is read, the parts enclosed in bars (“|”) will be entered verbatim into the symbol’s name. The “|” characters are not part of the symbol; they only serve to delimit the sequence of characters that must be entered “as is”. In order to maintain read-write invariance, symbols containing such sequences of special characters will be written between a pair of “|”

2Last-pair was a standard procedure in R^{3}RS.

'|x|  ⟹ x
(string->symbol "X")  ⟹ |X|
(symbol->string '|X|)  ⟹ "X"
'|a  b|  ⟹ |a  b|
'a|B|c  ⟹ |aBc|
(write '|FoO|)  ⟹ writes the string "|FoO|"
(display '|FoO|)  ⟹ writes the string "FoO"

Note: This notation has been introduced because R^{4}RS states that case must not be significant in symbols whereas the Tk toolkit is case significant (or more precisely thinks it runs over Tcl which is case significant). However, symbols containing the character “|” itself still can’t be read in.

(symbol? obj )  procedure

Returns #t if obj is a symbol, otherwise returns #f.

(symbol? 'foo)  ⟹ #t
(symbol? (car '(a b)))  ⟹ #t
(symbol? "bar")  ⟹ #f
(symbol? 'nil)  ⟹ #t
(symbol? '())  ⟹ #f
(symbol? #f)  ⟹ #f
(symbol? :key)  ⟹ #f

(symbol->string symbol)  procedure
(string->symbol string)  procedure

Identical to R^{4}RS.

(gensym)  procedure
(gensym prefix)  procedure

Gensym creates a new symbol. The print name of the generated symbol consists of a prefix (which defaults to "G") followed by the decimal representation of a number. If prefix is specified, it must be a string.

(gensym)
          ⟹ |G100|
(gensym "foo-")
          ⟹ foo-101

6.5 Numbers

The only numbers recognized by STk are integers (with arbitrary precision) and reals (implemented as C double floats).

(number? obj )  procedure

Returns #t if obj is a number, otherwise returns #f.

(complex? obj )  procedure

Returns the same result as number?. Note that complex numbers are not implemented.

(real? obj )  procedure

Returns #t if obj is a float number, otherwise returns #f.

(rational? obj )  procedure

Returns the same result as number?. Note that rational numbers are not implemented.

(integer? obj )  procedure

Returns #t if obj is an integer, otherwise returns #f. Note: The STk interpreter distinguishes

between integers which fit in a C long int (minus 8 bits) and integers of arbitrary length (aka “bignums”). This should be transparent to the user, though.

(exact? z)  procedure

(inexact? z)  procedure

In this implementation, integers (C long int or “bignums”) are exact numbers and floats are inexact.

(= z_{1} z_{2} z_{3} …)  procedure
(< x_{1} x_{2} x_{3} …)  procedure
(> x_{1} x_{2} x_{3} …)  procedure
(<= x_{1} x_{2} x_{3} …)  procedure
(>= x_{1} x_{2} x_{3} …)  procedure
(zero? z)  procedure

(positive? z)  procedure
(negative? z)  procedure

(odd? z)  procedure
(even? z)  procedure
(max x_{1} x_{2} …)  procedure
(min x_{1} x_{2} …)  procedure
(+ z_{1} …)  procedure
(* z_{1} …)  procedure
(- z_{1} z_{2})  procedure
(- z)  procedure
(- z_{1} z_{2} …)  procedure
(/ z_{1} z_{2})  procedure
(/ z)  procedure
(/ z_{1} z_{2} …)  procedure
(abs x)  procedure
(quotient n_{1} n_{2})  procedure
(remainder n_{1} n_{2})  procedure
(modulo n_{1} n_{2})  procedure

(gcd n_{1} …)  procedure
(lcm n_{1} …)  procedure

Identical to R^{4}RS.

(numerator q)  procedure
(denominator q)  procedure

Not implemented.

(floor x)  procedure
(ceiling x)  procedure

(truncate x)  procedure

(round x)  procedure

Identical to R^{4}RS.

(rationalize x y)  procedure

not yet implemented.

(exp z)  procedure
(log z)  procedure
(sin z)  procedure
(cos z)  procedure
(tan z)  procedure
(asin z)  procedure
(acos z)  procedure
(atan z)  procedure
(atan y x)  procedure
(sqrt z)  procedure
(expt z_{1} z_{2})  procedure

Identical to R^{4}RS.

(make-rectangular x_{1} x_{2})  procedure

(make-polar x_{1} x_{2})  procedure

(real-part z)  procedure
(imag-part z)  procedure
(magnitude z)  procedure

(angle z)  procedure

These procedures are not implemented since complex numbers are not defined.

(exact->inexact z)  procedure
(inexact->exact z)  procedure
(number->string number)  procedure

(number->string number radix)  procedure

(string->number string)  procedure

(string->number string radix)  procedure

Identical to R^{4}RS.

STk Reference Manual  17
name value alternate name name value alternate name
nul  000  null  bs  010  backspace
soh  001  ht  011  tab
stx  002  nl  012  newline
etx  003  vt  013
eot  004  np  014  page
enq 005  cr  015  return
ack  006  so  016
bel  007  bell  si  017
dle  020  can 030
dc1 021  em  031
dc2 022  sub 032
dc3 023  esc  033  escape
dc4 024  fs  034
nak 025  gs  035
syn 026  rs  036
etb  027  us  037
sp  040  space
del  177  delete

Table 1: Valid character names

6.6 Characters

Table 1 gives the list of allowed character names together with their ASCII equivalent expressed in octal.

(char? obj )  procedure
(char=? char_{1} char_{2})  procedure
(char<? char_{1} char_{2})  procedure
(char>? char_{1} char_{2})  procedure
(char<=? char_{1} char_{2})  procedure
(char>=? char_{1} char_{2})  procedure
(char-ci=? char_{1} char_{2})  procedure
(char-ci<? char_{1} char_{2})  procedure
(char-ci>? char_{1} char_{2})  procedure
(char-ci<=? char_{1} char_{2})  procedure
(char-ci>=? char_{1} char_{2})  procedure

(char-alphabetic? char)  procedure
(char-numeric? char)  procedure
(char-whitespace? char)  procedure
(char-upper-case? letter)  procedure
(char-lower-case? letter)  procedure
(char->integer char)  procedure
(integer->char n)  procedure
(char-upcase char)  procedure

Sequence  Character inserted
\b  Backspace
\e  Escape
\n  Newline
\t  Horizontal Tab
\n  Carriage Return
\0abc  ASCII character with octal value abc
\<newline> None (permits to enter a string on several lines)

                \<other>    <other>

Table 2: String escape sequences

(char-downcase char)  procedure

Identical to R^{4}RS.

6.7 Strings

STk string constants allow the insertion of arbitrary characters by encoding them as escape sequences, introduced by a backslash (\). The valid escape sequences are shown in Table 2. For instance, the string

    "ab\040c\nd\
    e"

is the string consisting of the characters #\a, #\b,#\space, #\c, #\newline, #\d and #\e.

(string? obj )  procedure

(make-string k)  procedure

(make-string k char)  procedure
(string char …)  procedure

(string-length string)  procedure

(string-ref string k)  procedure
(string-set! string k char)  procedure
(string=? string_{1} string_{2})  procedure
(string-ci=? string_{1} string_{2})  procedure
(string<? string_{1} string_{2})  procedure
(string>? string_{1} string_{2})  procedure
(string<=? string_{1} string_{2})  procedure
(string>=? string_{1} string_{2})  procedure
(string-ci<? string_{1} string_{2})  procedure
(string-ci>? string_{1} string_{2})  procedure
(string-ci<=? string_{1} string_{2})  procedure
(string-ci>=? string_{1} string_{2})  procedure
(substring string start end)  procedure
(string-append string …)  procedure
(string->list string)  procedure

(list->string chars)  procedure

(string-copy string)  procedure
(string-fill! string char)  procedure

Identical to R^{4}RS.

(string-find? string_{1} string_{2})  procedure

Returns #t if string_{1} appears somewhere in string_{2}; otherwise returns #f.

(string-index string_{1} string_{2})  procedure

Returns the index of where string_{1} is a substring of string_{2} if it exists; returns #f otherwise.

    (string-index "ca" "abracadabra")
              ⟹ 4
    (string-index "ba" "abracadabra")
              ⟹ #f

(split-string string)  procedure
(split-string string delimiters)  procedure

This function parses string_{1} and returns a list of tokens ended by a character of the delimiters_{1} string. If delimiters_{1} is omitted, it defaults to a string containing a space, a tabulation and a newline characters.

    (split-string "/usr/local/bin" "/") ⟹ ("usr" "local" "bin")
    (split-string "once   upon a time") ⟹ ("once" "upon" "a" "time")

(string-lower string)  procedure

Returns a string in which all upper case letters of string have been replaced by their lower case equivalent.

(string-upper string)  procedure

Returns a string in which all lower case letters of string have been replaced by their upper case equivalent.

6.8 Vectors

(vector? obj )  procedure

(make-vector k)  procedure
(make-vector k fill)  procedure

(vector obj …)  procedure

(vector-length vector)  procedure

(vector-ref vector k)  procedure
(vector-set! vector k obj )  procedure

(vector->list vector)  procedure

(list->vector list)  procedure

(vector-fill! vector fill)  procedure

Identical to R^{4}RS.

(vector-copy vector)  procedure

returns a copy of vector.

(vector-resize vector size)  procedure

vector-resize physically changes the size of vector. If size is greater than the old vector size, the contents of the newly allocated cells are undefined.

6.9 Control features

(procedure? obj )  procedure

(apply proc args)  procedure
(apply proc arg_{1} … args)  procedure
(map proc list_{1} list_{2} …)  procedure
(for-each proc list_{1} list_{2} …)  procedure
(force promise)  procedure

Identical to R^{4}RS.

(call-with-current-continuation proc)  procedure

(call/cc proc)  procedure

Call/cc is a shorter name for call-with-current-continuation.

(closure? obj )  procedure

returns #t if obj is a procedure created by evaluating a lambda expression, otherwise returns

#f.

(primitive? obj )  procedure

returns #t if obj is a procedure and is not a closure, otherwise returns #f.

(promise? obj )  procedure

returns #t if obj is an object returned by the application of delay, otherwise returns #f.

(continuation? obj )  procedure

returns #t if obj is a continuation obtained by call/cc, otherwise returns #f.

(dynamic-wind ⟨thunk_{1}⟩ ⟨thunk_{2}⟩ ⟨thunk_{3}⟩)  procedure

⟨Thunk_{1}⟩, ⟨thunk_{2}⟩ and ⟨thunk_{3}⟩ are called in order. The result of dynamic-wind is the value returned by ⟨thunk_{2}⟩. If ⟨thunk_{2}⟩ escapes from its continuation during evaluation (by calling a continuation obtained by call/cc or on error), ⟨thunk_{3}⟩ is called. If ⟨thunk_{2}⟩ is later reentered, ⟨thunk_{1}⟩ is called.

(catch ⟨expression_{1}⟩ ⟨expression_{2}⟩ …)  syntax

The ⟨expression⟩s are evaluated from left to right. If an error occurs, evaluation of the ⟨expression⟩s is aborted, and #t is returned to catch’s caller. If evaluation finishes without an error, catch returns #f.

    (let* ((x 0)
           (y (catch
                 (set! x 1)
                 (/ 0) ; causes a "division by 0" error
                 (set! x 2))))
      (cons x y))
              ⟹ (1 . #t)

(procedure-body ⟨procedure⟩)  procedure

returns the body of ⟨procedure⟩. If ⟨procedure⟩ is not a closure, procedure-body returns

#f.
    (define (f a b)
       (+ a (* b 2)))
    (procedure-body f)  ⟹ (lambda (a b)
                                      (+ a (* b 2)))
    (procedure-body car)  ⟹ #f

6.10 Input and output

The R^{4}RS states that ports represent input and output devices. However, it defines only ports which are attached to files. In STk, ports can also be attached to strings or to a external command input or output. String ports are similar to file ports, except that characters are read from (or written to) a string rather than a file. External command input or output ports are implemented with Unix pipes and are called pipe ports. A pipe port is created by specifying the command to execute prefixed with the string "| ". Specification of a pipe port can occur everywhere a file name is needed.

(call-with-input-file string proc)  procedure
(call-with-output-file string proc)  procedure

Note: if string starts with the two characters "| ", these procedures return a pipe port. Consequently, it is not possible to open a file whose name starts with those two characters.

(call-with-input-string string proc)  procedure

behaves exactly as call-with-input-file except that the port passed to proc is the string port obtained from string.

    (call-with-input-string "123 456" (lambda (x) (read x)))
              ⟹ 123

(call-with-output-string proc)  procedure

Proc should be a procedure of one argument. Call-with-output-string calls proc with a freshly opened output string port. The result of this procedure is a string containing all the text that has been written on the string port.

    (call-with-output-string
       (lambda (x) (write 123 x) (display "Hello" x)))
              ⟹ "123Hello"

(input-port? obj )  procedure
(output-port? obj )  procedure

Identical to R^{4}RS.

(input-string-port? obj )  procedure
(output-string-port? obj )  procedure

Returns #t if obj is either an input or an output string port, otherwise returns #f.

(current-input-port)  procedure
(current-output-port)  procedure

Identical to R^{4}RS.

(current-error-port)  procedure

Returns the current default error port.

(with-input-from-file string thunk)  procedure
(with-output-to-file string thunk)  procedure

Identical to R^{4}RS. The following example uses a pipe port opened for reading. It permits to read all the lines produced by an external ls command (i.e. the ouput of the ls command is redirected to the Scheme pipe port).

    (with-input-from-file "| ls -ls"
      (lambda ()
        (do ((l (read-line) (read-line)))
            ((eof-object? l))
          (display l)
          (newline))))

Hereafter is another example of Unix command redirection. This time, it is the standard input of the Unix command which is redirected.

    (with-output-to-file "| mail root"
      (lambda()
        (format #t "A simple mail sent from STk\n")))

(with-input-from-string string thunk)  procedure

A string port is opened for input from string. Current-input-port is set to the port and thunk is called. When thunk returns, the previous default input port is restored. With-input-from-string returns the value yielded by thunk.

    (with-input-from-string "123 456" (lambda () (read)))
              ⟹ 123

(with-output-to-string thunk)  procedure

A string port is opened for output. Current-output-port is set to it and thunk is called. When the thunk returns, the previous default output port is restored. With-output-to-string returns the string containing all the text written on the string port.

    (with-output-to-string (lambda () (write 123) (write "Hello")))
              ⟹ "123Hello"

(open-input-file filename)  procedure
(open-output-file filename)  procedure

Identical to R^{4}RS.

Note: if f ilename starts with the string "| ", these procedure return a pipe port. Consequently, it is not possible to open a file whose name starts with those two characters.

(open-input-string string)  procedure

Returns an input string port capable of delivering characters from string.

(open-output-string)  procedure

Returns an output string port capable of receiving and collecting characters.

(get-output-string port)  procedure

Returns a string containing all the text that has been written on the output string port.

    (let ((p (open-output-string)))
      (display "Hello, world" p)
      (get-output-string p))
              ⟹ "Hello, world"

(close-input-port port)  procedure
(close-output-port port)  procedure

Identical to R^{4}RS.

(read)  procedure
(read port)  procedure

(read-char)  procedure

(read-char port)  procedure

(peek-char)  procedure

(peek-char port)  procedure

(char-ready?)  procedure
(char-ready? port)  procedure

Identical to R^{4}RS.

(read-line)  procedure

(read-line port)  procedure

Reads the next line available from the input port port and returns it as a string. The terminating newline is not included in the string. If no more characters are available, an end of file object is returned. Port may be omitted, in which case it defaults to the value returned

by current-input-port.

(write obj )  procedure
(write obj port)  procedure
(display obj )  procedure
(display obj port)  procedure

(newline)  procedure

(newline port)  procedure
(write-char char)  procedure
(write-char char port)  procedure

Identical to R^{4}RS.

(format port string obj_{1} obj_{2} …)  procedure

Writes the obj s to the given port, according to the format string string. String is written literally, except for the following sequences:

• ~a or ~A is replaced by the printed representation of the next obj .

• ~s or ~S is replaced by the “slashified” printed representation of the next obj .

• ~~ is replaced by a single tilde.

• ~% is replaced by a newline

Port can be a boolean, a port or a string port. If port is #t, output goes to the current output port; if port is #f, the output is returned as a string. Otherwise, the output is printed on the specified port.

    (format #f "A test.")
              ⟹ "A test."
    (format #f "A ~a." "test")
              ⟹ "A test."
    (format #f "A ~s." "test")
              ⟹ "A \"test\"."

(get-output-string port)  procedure

Returns the string associated with the output string port.

    (let ((p (open-output-string)))
      (display "Hello, world" p)
      (get-output-string p))
              ⟹ "Hello, world"

(flush)  procedure
(flush port)  procedure

Flushes the buffer associated with the given port. The port argument may be omitted, in which case it defaults to the value returned by current-output-port.

(when-port-readable port handler)  procedure
(when-port-readable port)  procedure

When port is ready for reading, handler, which must be a thunk, is called leaving the current evaluation suspended. When handler execution is terminated, normal evaluation is resumed at its suspension point. If the special value #f is provided as handler, the current handler for port is deleted. If a handler is provided, the value returned by when-port-readable is undefined. Otherwise, it returns the handler currently associated to port. The example below shows a simple usage of the when-port-readable procedure: the command cmd is run with its output redirected in a pipe associated to the p Scheme port.

    (define p (open-input-file "| cmd"))
    (when-port-readable p
                (lambda()
                  (let (( l (read-line p)))
                    (if (eof-object? l)
                        (begin
                          ;; delete handler
                          (when-port-readable p #f)
                          ;; and close port
                          (close-input-port p))
                        (format #t "Line read: ~A\n" l)))))

(when-port-writable port handler)  procedure
(when-port-writable port)  procedure

When port is ready for writing, handler, which must be a thunk, is called leaving the current evaluation suspended. When handler execution is terminated, normal evaluation is resumed at its suspension point. If the special value #f is provided as handler, the current handler for port is deleted. If a handler is provided, the value returned by when-port-writable is undefined. Otherwise, it returns the handler currently associated to port.

(load filename)  procedure
(load filename module)  procedure

The first form is identical to R^{4}RS. The second one loads the content of filename in the given

module environment. Note: The load primitive has been extended to allow loading of object files, though this is not implemented on all systems. This extension uses dynamic loading on systems which support it ^{3}. See [4] for more details.

(try-load filename)  procedure
(try-load filename module)  procedure

Tries to load the file named filename. If filename exists and is readable, it is loaded, and try-load returns #t. Otherwise, the result of the call is #f. The second form of try-load tries to load the content of filename in the given module environment.

(autoload filename ⟨symbol_{1}⟩ ⟨symbol_{2}⟩ …)  syntax

Defines ⟨symbol⟩s as autoload symbols associated to file filename. First evaluation of an autoload symbol will cause the loading of its associated file in the module environment in which the autoload was done. Filename must provide a definition for the symbol which lead to its loading, otherwise an error is signaled.

(autoload? symbol module)  procedure

Returns #t if symbol is an autoload symbol in module environment ; returns #f otherwise.

(require string)  procedure
(provide string)  procedure
(provided? string)  procedure

Require loads the file whose name is string if it was not previously “provided”.Provide permits to store string in the list of already provided files. Providing a file permits to avoid subsequent loads of this file. Provided? returns #t if string was already provided; it returns #f otherwise.

(transcript-on filename)  procedure

(transcript-off)  procedure

Not implemented.

3Current version (3.99) allows dynamic loading only on some platforms: SunOs 4.1.x, SunOs 5.x, NetBSD 1.0, Linux 2.0, HPUX, Irix 5.3

(open-file filename mode)  procedure

Opens the file whose name is filename with the specified mode. Mode must be “r” to open for reading or “w” to open for writing. If the file can be opened, open-file returns the port associated with the given file, otherwise it returns #f. Here again, the “magic” string "| `` permit to open a pipe port.

(close-port port)  procedure

Closes port. If port denotes a string port, further reading or writing on this port is disallowed.

(transcript-on filename)  procedure

(transcript-off)  procedure

Not implemented.

(port->string port)  procedure

(port->list reader port)  procedure

(port->string-list port)  procedure
(port->sexp-list port)  procedure

Those procedures are utility for generally parsing input streams. Their specification has been stolen from scsh. Port->string reads the input port until eof, then returns the accumulated string.

    (port->string (open-input-file "| (echo AAA; echo BBB)"))
                                ⟹ "AAA\nBBB\n"
    (define exec
            (lambda (command)
               (call-with-input-file
                  (string-append "| " command) port->string)))
    (exec "ls -l")  ⟹ a string which contains the result of "ls -l"

Port->list uses the reader function to repeatedly read objects from port. Thes objects are accumulated in a list which is returned upon eof.

    (port->list read-line (open-input-file "| (echo AAA; echo BBB)"))
                                ⟹ ("AAA" "BBB")

Port->string-list reads the input port line by line until eof, then returns the accumulated list of lines. This procedure is defined as

    (define port->string-list (lambda (p)(port->list read-line p)))

Port->sexp-list repeatedly reads data from the port until eof, then returns the accumulated list of items. This procedure is defined as

    (define port->sexp-list (lambda (p) (port->list read p)))

For instance, the following expression gives the list of users currently connected on the machine running the STk interpreter.

    (port->sexp-list (open-input-file "| users"))

6.11 Keywords

Keywords are symbolic constants which evaluate to themselves. A keyword must begin with a colon.

(keyword? obj )  procedure

Returns #t if obj is a keyword, otherwise returns #f.

(make-keyword obj )  procedure

Builds a keyword from the given obj . obj must be a symbol or a string. A colon is automatically prepended.

    (make-keyword "test")
              ⟹ :test
    (make-keyword 'test)
              ⟹ :test
    (make-keyword ":hello")
              ⟹ ::hello

(keyword->string keyword)  procedure

Returns the name of keyword as a string. The leading colon is included in the result.

    (keyword->string :test)
              ⟹ ":test"

(get-keyword keyword list)  procedure
(get-keyword keyword list default)  procedure

List must be a list of keywords and their respective values. Get-keyword scans the list and returns the value associated with the given keyword. If the keyword does not appear in an odd position in list, the specified default is returned, or an error is raised if no default was specified.

    (get-keyword :one '(:one 1 :two 2))
              ⟹ 1
    (get-keyword :four '(:one 1 :two 2) #f)
              ⟹ #f
    (get-keyword :four '(:one 1 :two 2))
              ⟹ error

6.12 Tk commands

As we mentioned in the introduction, STk can easily communicate with the Tk toolkit. All the commands defined by the Tk toolkit are visible as Tk-commands, a basic type recognized by the interpreter. Tk-commands can be called like regular scheme procedures, serving as an entry point into the Tk library.

Note: Some Tk-commands can dynamically create other Tk-commands. For instance, execution of the expression

    (label '.lab)

will create a new Tk-command called “.lab”. This new object, which was created by a primitive Tk-command, will be called a widget. Note: When a new widget is created, it captures its creation environment. This permits to have bindings which access variables in the scope of the widget creation call (see 6.17).

(tk-command? obj )  procedure

Returns #t if obj is a Tk-command, otherwise returns #f.

    (tk-command? label)
              ⟹ #t
    (begin (label '.lab) (tk-command? .lab))
              ⟹ #t
    (tk-command? 12)
              ⟹ #f

(widget? obj )  procedure

Returns #t if obj is a widget, otherwise returns #f. A widget is a Tk-command created by a

primitive Tk-command such as button, label, menu, etc.
    (widget? label)
              ⟹ #f
    (begin (label '.lab) (widget? .lab))
              ⟹ #t
    (widget? 12)
              ⟹ #f

(widget->string widget)  procedure

Returns the widget name of widget as a string.

    (begin (label '.lab) (widget->string .lab))
              ⟹ ".lab"

(string->widget str)  procedure

Returns the widget whose name is str if it exists; otherwise returns #f.

    (begin (label '.lab) (string->widget ".lab"))
              ⟹ the Tk-command named ".lab"

(widget-name widget)  procedure

Returns the widget name of widget as a symbol.

    (begin (label '.lab) (widget->name .lab))
              ⟹ .lab

(set-widget-data! widget expr)  procedure

Set-widget-data! associates arbitrary data with a widget. The system makes no assumptions about the type of expr; the data is for programmer convenience only. As shown below, it could be used as a kind of property list for widgets.

(get-widget-data widget)  procedure

Returns the data previously associated with widget if it exists; otherwise returns #f.

    (begin
       (set-widget-data! .w '(:mapped #t :geometry "10x50"))
       (get-keyword :mapped (get-widget-data .w)))
              ⟹ #t

6.13 Modules

STk modules can be used to organize a program into separate environments (orname spaces). Modules provide a clean way to organize and enforce the barriers between the components of a program. STk provides a simple module system which is largely inspired from the one of Tung and Dybvig exposed in [5]. As their modules system, STk modules are defined to be easily used in an interactive environment.

(define-module name ⟨body⟩)  syntax

Define-module evaluates the expressions which are in ⟨body⟩ in the environment of the module name. name must be a valid symbol. If this symbol has not already been used to define a module, a new module, named name, is created. Otherwise, ⟨body⟩ is evaluated in the environment of the (old) module name^{4}. Definitions done in a module are local to the module and do not interact with the definitions of other modules. Consider the following definitions,

4In fact define-module on a given name defines a new module only the first time it is invoked on this name.

By this way, inteactively reloading a module does not define a new entity, and the other modules which use it are not altered.

      (define-module M1
         (define a 1))
      (define-module M2
        (define a 2)
        (define b (* 2 x)))

Here, two modules are defined and they both bind the symbol a to a value. However, since a has been defined in two distincts modules they denote two different locations. The “STk” module, which is predefined, is a special module which contains all the global variables of a R^{4}RS program. A symbol defined in the STk module, if not hidden by a local definition, is always visible from inside a module. So, in the previous exemple, the x symbol refers the x symbol defined in the STk module. The result of define-module is undefined.

(find-module name)  procedure

(find-module name default)  procedure

STk modules are first class objects and find-module returns the module associated to name if it exists. If there is no module associated to name, an error is signaled if no default is provided, otherwise find-module returns default.

(module? object)  procedure

Returns #t if object is a module and #f otherwise.

      (module? (find-module 'STk))
              ⟹ #t
      (module? 'STk)
              ⟹ #f
      (module? 1)
              ⟹ #f

(export ⟨symbol_{1}⟩ ⟨symbol_{2}⟩…)  syntax

Specifies the symbols which are exported (i.e. visible) outside the current module. By default, symbols defined in a module are not visible outside this module, excepted the symbols which appear in an export clause. If several export clauses appear in a module, the set of exported symbols is determined by unioning symbols exported in all the export clauses. The result of export is undefined.

(import ⟨module_{1}⟩ ⟨module_{2}⟩…)  syntax

Specifies the modules which are imported by the current module. Importing a module makes the symbols it exports visible to the importer, if not hidden by local definitions. When a symbol is exported by several of the imported modules, the location denoted by this symbol in the importer module correspond to the one of the first module in the list (⟨module_{1}⟩ ⟨module_{2}⟩…) which export it. If several import clauses appear in a module, the set of imported modules is determined by appending the various list of modules in their apparition order.

(define-module M1
(export a b)
(define a 'M1-a)
(define b 'M1-b))
(define-module M2
(export b c)
(define b 'M2-b)
(define c 'M2-c))
(define-module M3
(import M1 M2)
(display (list a b c)))
        ⟹ displays (m1-a m1-b m2-c)

Note: There is no kind of transitivity in module importations: when the module C imports the module B which an importer of A, the symbols of A are not visible from C, except by explicitly importing the A module from C. Note: The module STk, which contains the global variables is always implicitly imported from a module.Furthermore, this module is always placed at the end of the list of imported modules.

(export-symbol symbol module)  procedure

Exports symbol from module. This procedure can be useful, when debugging a program, to make visible a given symbol without reloading or redefining the module where this symbol was defined.

(export-all-symbols)  procedure

Exports all the symbols of current module . If symbols are added to the current module after the call to export-all-symbols, they are automatically exported.

Note: The STk module export all the symbols which are defined in it (i.e. global variables are visible, if not hidden, from all the modules of a program.

(with-module name ⟨expr_{1}⟩ ⟨expr_{2}⟩ …)  syntax

Evaluates the expressions of ⟨expr_{1}⟩ ⟨expr_{2}⟩ … in the environment of module name. Module name must have been created previously by a define-module. The result of with-module is the result of the evaluation of the last ⟨expr⟩.

(define-module M
  (define a 1)
  (define b 2))
(with-module M
   (+ a b))
        ⟹ 3

(current-module)  procedure

Returns the current-module.

      (define-module M
         ...)
       (with-module M
         (cons (eq? (current-module) (find-module 'M))
               (eq? (current-module) (find-module 'STk))))
              ⟹ (#t . #f)

(select-module name)  syntax

Evaluates the expressions which follows in module name environment. Module name must have been created previously by a define-module. The result of select-module is undefined. Select-module is particularly useful when debugging since it allows to place toplevel evaluation in a particular module. The following transcript shows an usage of select-module^{5}:

    STk> (define foo 1)
    STk> (define-module bar
            (define foo 2))
    STk> foo
    1
    STk> (select-module bar)
    bar> foo
    2
    bar> (select-module STk)
    STk>

(module-name module)  procedure

Returns the name (a symbol) associated to a module.

(module-imports module)  procedure

Returns the list modules that module imports.

(module-exports module)  procedure

Returns the list of symbols exported by module.

(module-symbols module)  procedure

Returns the list symbols that ere defined in module.

5This transcript uses the default value for the function repl-display-prompt (see page 72) which displays the name of the current module in the prompt.

6.14 Environments

Environments are first class objects in STk. The following primitives are defined on environments.

(environment? obj )  procedure

Returns #t if obj is an environment, otherwise returns #f.

(the-environment)  procedure

Returns the current environment.

(global-environment)  procedure

Returns the “global” environment (i.e. the toplevel environment).

(parent-environment env)  procedure

Returns the parent environment of env. If env is the “global” environment (i.e. the toplevel environment), parent-environment returns #f.

(environment->list environment)  procedure

Returns a list of a-lists, representing the bindings in environment. Each a-list describes one level of bindings, with the innermost level coming first.

    (define E (let ((a 1) (b 2))
                (let ((c 3))
                  (the-environment))))
    (car (environment->list E)) ⟹ ((c . 3))
    (cadr (environment->list E))⟹ ((b . 2) (a . 1))

(procedure-environment procedure)  procedure

Returns the environment associated with procedure. Procedure-environment returns #f if procedure is not a closure.

    (define foo (let ((a 1)) (lambda () a)))
    (car (environment->list
             (procedure-environment foo)))
                                ⟹ ((a . 1))

(module-environment module)  procedure

Returns the environment associated to the module module.

    (define-module M
       (define a 1))
    (car (environment->list
             (module-environment (find-module 'M))))
                                ⟹ ((a . 1))

(symbol-bound? symbol)  procedure

(symbol-bound? symbol environment)  procedure

Returns #t if symbol has a value in the given environment, otherwise returns #f. Environment may be omitted, in which case it defaults to the global environment.

6.15 Macros

STk provides low level macros.

Note: STk macros are not the sort of macros defined in the appendix of R^{4}RS, but rather the macros one can find in most of Lisp dialects.

(macro ⟨formals⟩ ⟨body⟩)  syntax

Macro permits to create a macro. When a macro is called, the whole form (i.e. the macro itself and its parameters) is passed to the macro body. Binding association is done in the environment of the call. The result of the binding association is called the macro-expansion. The result of the macro call is the result of the evaluation of the macro expansion in the call environment.

    (define foo (macro f ` (quote ,f)))
    (foo 1 2 3)  ⟹ (foo 1 2 3)
    (define 1+ (macro form (list + (cadr form) 1)))
    (let ((x 1)) (1+ x))  ⟹ 2

(macro? obj )  procedure

Returns #t if obj is a macro, otherwise returns #f.

(macro-expand-1 form)  procedure
(macro-expand form)  procedure

Macro-expand-1 returns the macro expansion of form if it is a macro call, otherwise form is returned unchanged. Macro-expand is similar to macro-expand-1, but repeately expand form until it is no longer a macro call.

    (define 1- (macro form ` (-  ,(cadr form) 1)))
    (define -- (macro form ` (1- ,(cadr form))))
    (macro-expand-1 '(1- 10)) ⟹ (- 10 1)
    (macro-expand   '(1- 10)) ⟹ (- 10 1)
    (macro-expand-1 '(-- 10)) ⟹ (1- 10)
    (macro-expand   '(-- 10)) ⟹ (- 10 1)

(macro-expand form)  procedure

Returns the macro expansion of form if it is a macro call, otherwise form is returned unchanged. Macro expansion continue until, the form obtained is

(define 1- (macro form (list '- (cadr form) 1)))
(macro-expand '(1- 10))  ⟹ (- 10 1)

(macro-body macro)  procedure

Returns the body of macro

(macro-body 1+)
                            ⟹ (macro form (list + (cadr form) 1))

(define-macro (⟨name⟩ ⟨formals⟩) ⟨body⟩)  macro

Define-macro is a macro which permits to define a macro more easily than with the macro form. It is similar to the defmacro of Common Lisp [6].

(define-macro (incr x) ` (set! ,x (+ ,x 1)))
(let ((a 1)) (incr a) a)  ⟹ 2
(define-macro (when test . body)
  ` (if ,test ,@(if (null? (cdr body)) body ` ((begin ,@body)))))
(macro-expand '(when a b)) ⟹ (if a b)
(macro-expand '(when a b c d))
                            ⟹ (if a (begin b c d))

Note: Calls to macros defined by define-macro are physically replaced by their macro-expansion if the variable *debug* is #f (i.e. their body is “in-lined” in the macro call). To avoid this feature, and to ease debugging, you have to set this variable to #t. (See also 6.25).

6.16 System procedures

This section lists a set of procedures which permits to access some system internals.

(expand-file-name string)  procedure

Expand-file-name expands the filename given in string to an absolute path. This function understands the tilde convention for filenames.

;; Current directory is /users/eg/STk
(expand-file-name "..")
          ⟹ "/users/eg"
(expand-file-name "~root/bin)
          ⟹ "/bin"
(expand-file-name "~/STk)"
          ⟹ "/users/eg/STk"

(canonical-path path)  procedure

Expands all symbolic links in path and returns its canonicalized absolute pathname. The resulting path do not have symbolic links. If path doesn’t designate a valid pathname, canonical-path returns #f.

(dirname string)  procedure

Returns a string containing all but the last component of the path name given in string.

    (dirname "/a/b/c.stk")
              ⟹ "/a/b"

(basename string)  procedure

Returns a string containing the last component of the path name given in string.

    (basname "/a/b/c.stk")
              ⟹ "c.stk"

(decompose-file-name string)  procedure

Returns an “exploded” list of the path name components given in string. The first element in the list denotes if the given string is an absolute path or a relative one, being "/" or "." respectively. Each component of this list is a string.

    (decompose-file-name "/a/b/c.stk")
              ⟹ ("/" "a" "b" "c.stk")
    (decompose-file-name "a/b/c.stk")
              ⟹ ("." "a" "b" "c.stk")

(file-is-directory? string)  procedure
(file-is-regular? string)  procedure
(file-is-readable? string)  procedure
(file-is-writable? string)  procedure
(file-is-executable? string)  procedure

(file-exists? string)  procedure

Returns #t if the predicate is true for the path name given in string; returns #f otherwise (or if string denotes a file which does not exist).

(glob pattern_{1} pattern_{2} …)  procedure

The code for glob is taken from the Tcl library. It performs file name “globbing” in a fashion similar to the csh shell. Glob returns a list of the filenames that match at least one of the pattern arguments. The pattern arguments may contain the following special characters:

• ? Matches any single character.

• * Matches any sequence of zero or more characters.

• [chars] Matches any single character in chars. If chars contains a sequence of the form a-b then any character between a and b (inclusive) will match.

• \x Matches the character x.

• {a,b,...} Matches any of the strings a, b, etc.

As with csh, a “.” at the beginning of a file’s name or just after a “/” must be matched explicitly or with a {} construct. In addition, all “/” characters must be matched explicitly. If the first character in a pattern is “~” then it refers to the home directory of the user whose name follows the “~”. If the “~” is followed immediately by “/” then the value of the environment variable HOME is used. Glob differs from csh globbing in two ways. First, it does not sort its result list (use the sort procedure if you want the list sorted). Second, glob only returns the names of files that actually exist; in csh no check for existence is made unless a pattern contains a ?, *, or []construct.

(remove-file string)  procedure

Removes the file whose path name is given in string. The result of remove-file is undefined.

(rename-file string_{1} string_{2})  procedure

Renames the file whose path-name is contained in string_{1} in the path name given by string_{2}. The result of rename-file is undefined.

(temporary-file-name string)  procedure

Generates a unique temporary file name. The value returned by temporary-file-name is the newly generated name of #f if a unique name cannot be generated.

(getcwd)  procedure

Getcwd returns a string containing the current working directory.

(chdir string)  procedure

Chdir changes the current directory to the directory given in string.

(getpid string)  procedure

Returns the system process number of the current STk interpreter (i.e. the Unix pid). Result is an integer.

(system string)  procedure
(! string)  procedure

Sends the given string to the system shell /bin/sh. The result of system is the integer status code the shell returns.

(exec string)  procedure

Executes the command contained in string and redirects its output in a string. This string constitutes the result of exec.

(getenv string)  procedure

Looks for the environment variable named string and returns its value as a string, if it exists. Otherwise, getenv returns #f.

    (getenv "SHELL")
              ⟹ "/bin/zsh"

(setenv! var value)  procedure

Sets the environment variable var to value. Var and value must be strings. The result of setenv! is undefined.

    (getenv "SHELL")
              ⟹ "/bin/zsh"

6.17 Addresses

An address is a Scheme object which contains a reference to another Scheme object. This type can be viewed as a kind of pointer to a Scheme object. Addresses, even though they are very dangerous, have been introduced in STk so that objects that have no “readable” external representation can still be transformed into strings and back without loss of information. Adresses were useful with pre-3.0 version of STk; their usage is now stongly discouraged, unless you know what you do. In particular, an address can designate an object at a time and another one later (i.e. after the garbage collector has marked the zone as free). Addresses are printed with a special syntax: #pNNN, where NNN is an hexadecimal value. Reading this value back yields the original object whose location is NNN.

(address-of obj )  procedure

Returns the address of obj.

(address? obj )  procedure

Returns #t if obj is an address; returns #f otherwise.

6.18 Signals

STk allows the use to associate handlers to signals. Signal handlers for a given signal can even be chained in a list. When a signal occurs, the first signal of the list is executed. Unless this signal yields the symbol break the next signal of the list is evaluated. When a signal handler is called, the integer value of this signal is passed to it as (the only) parameter. The following POXIX.1 constants for signal numbers are defined: SIGABRT, SIGALRM, SIGFPE,

SIGHUP, SIGILL, SIGINT, SIGKILL, SIGPIPE, SIGQUIT, SIGSEGV, SIGTERM, SIGUSR1, SIGUSR2,
SIGCHLD, SIGCONT, SIGSTOP, SIGTSTP, SIGTTIN, SIGTTOU. Moreover, the following constants,

which are often available on most systems are also defined^{6}: SIGTRAP, SIGIOT, SIGEMT,

6Some of these constants may be undefined if they are not supported by your system

SIGBUS, SIGSYS, SIGURG, SIGCLD, SIGIO, SIGPOLL, SIGXCPU, SIGXFSZ, SIGVTALRM, SIGPROF,
SIGWINCH, SIGLOST.

See your Unix documentation for the exact meaning of each constant or [7]. Use symbolic constants rather than their numeric value if you plan to port your program on another system. A special signal, managed by the interpreter, is also defined: SIGHADGC. This signal is raised when the garbage collector phase terminates. When the interpreter starts running, all signals are sets to their default value, excepted SIGINT (generally bound to Control-C) which is handled specially.

(set-signal-handler! sig handler)  procedure

Replace the handler for signal sig with handler. Handler can be

- #t to reset the signal handler for sig to the default system handler.

- #f to completly ignore sig (Note that Posix.1 states that SIGKILL and SIGSTOP cannot be caught or ignored).

- a one parameter procedure.

This procedure returns the new handler, or (length 1) handler list, associated to sig.

    (let* ((x       #f)
           (handler (lambda (i) (set! x #t))))
      (set-signal-handler! |SIGHADGC| handler)
      (gc)
      x)
              ⟹ #t

(add-signal-handler! sig handler)  procedure

Adds handler to the list of handlers for signal sig. If the old signal handler is a boolean, this procedure is equivalent to set-signal-handler!. Otherwise, the new handler is added in front of the previous list of handler. This procedure returns the new handler, or handler list, associated to sig.

    (let* ((x        '())
           (handler1 (lambda (i) (set! x (cons 1 x))))
           (handler2 (lambda (i) (set! x (cons 2 x)))))
      (add-signal-handler! |SIGHADGC| handler1)
      (add-signal-handler! |SIGHADGC| handler2)
      (gc)
      x)
              ⟹ (1 2)

    (let* ((x        '())
           (handler1 (lambda (i) (set! x (cons 1 x))))
           (handler2 (lambda (i) (set! x (cons 2 x)) 'break)))

      (add-signal-handler! |SIGHADGC| handler1)
      (add-signal-handler! |SIGHADGC| handler2)
      (gc)
      x)
              ⟹ (2)

(get-signal-handlers)  procedure
(get-signal-handlers sig)  procedure

Returns the handlers, or the list of handlers, associated to the signal sig. If sig is omitted, get-signal-handlers returns a vector of all the signal handlers currently in effect.

6.19 Hash tables

A hash table consists of zero or more entries, each consisting of a key and a value. Given the key for an entry, the hashing function can very quickly locate the entry, and hence the corresponding value. There may be at most one entry in a hash table with a particular key, but many entries may have the same value. STk hash tables grow gracefully as the number of entries increases, so that there are always less than three entries per hash bucket, on average. This allows for fast lookups regardless of the number of entries in a table.

Note: Hash table manipulation procedures are built upon the efficient Tcl hash table package.

(make-hash-table)  procedure

(make-hash-table comparison)  procedure
(make-hash-table comparison hash)  procedure

Make-hash-table admits three different forms. The most general form admit two arguments. The first argument is a comparison function which determine how keys are compared; the second argument is a function which computes a hash code for an object and returns the hash code as a non negative integer. Objets with the same hash code are stored in an A-list registered in the bucket corresponding to the key. If omitted,

• hash defaults to the hash-table-hash procedure.

• comparison defaults to the eq? procedure

Consequently,

    (define h (make-hash-table))

is equivalent to

    (define h (make-hash-table eq? hash-table-hash))

Another interesting example is

    (define h (make-hash-table string-ci=? string-length))

which defines a new hash table which uses string-ci=? for comparing keys. Here, we use the string-length as a (very simple) hashing function. Of course, a function which gives a key depending of the characters composing the string gives a better repartition and should probably enhance performances. For instance, the following call to make-hash-table should return a more efficient, even if not perfect, hash table:

    (make-hash-table
        string-ci=?
        (lambda (s)
          (let ((len (string-length s)))
            (do ((h 0)  (i 0 (+ i 1)))
                ((= i len) h)
              (set! h (+ h (char->integer
                             (char-downcase (string-ref s i)))))))))

Note: Hash tables with a comparison function equal to eq? or string=? are handled in an more efficient way (in fact, they don’t use the hash-table-hash fucntion to speed up hash table retrievals).

(hash-table? obj )  procedure

Returns #t if obj is a hash table, returns #f otherwise.

(hash-table-hash obj )  procedure

hash-table-hash computes a hash code for an object and returns the hash code as a non negative integer. A property of hash-table-hash is that

     (equal?  x y) implies (equal?  (hash-table-hash x) (hash-table-hash y)

as the the Common Lisp sxhash function from which this procedure is modeled.

(hash-table-put! hash key value)  procedure

Hash-table-put! enters an association between key and value in the hash table. The value returned by hash-table-put! is undefined.

(hash-table-get hash key)  procedure

(hash-table-get hash key default)  procedure

Hash-table-get returns the value associated with key in the given hash table. If no value has been associated with key in hash, the specified default is returned if given; otherwise an error is raised.

    (define h1 (make-hash-table))
    (hash-table-put! h1 'foo (list 1 2 3))
    (hash-table-get  h1 'foo)
              ⟹ (1 2 3)
    (hash-table-get  h1 'bar 'absent)
              ⟹ absent

    (hash-table-get  h1 'bar)
              ⟹ error
    (hash-table-put! h1 '(a b c) 'present)
    (hash-table-get  h1 '(a b c) 'absent)
              ⟹ 'absent
    (define h2 (make-hash-table equal?))
    (hash-table-put! h2 '(a b c) 'present)
    (hash-table-get  h2 '(a b c))
              ⟹ 'present

(hash-table-remove! hash key)  procedure

hash must be a hash table containing an entry for key. Hash-table-remove! deletes the entry for key in hash, if it exists. Result of Hash-table-remove! is unspecified.

    (define h (make-hash-table))
    (hash-table-put! h 'foo (list 1 2 3))
    (hash-table-get h 'foo)
              ⟹ (1 2 3)
    (hash-table-remove! h 'foo)
    (hash-table-get h 'foo 'absent)
              ⟹ absent

(hash-table-for-each hash proc)  procedure

Proc must be a procedure taking two arguments. Hash-table-for-each calls proc on each key/value association in hash, with the key as the first argument and the value as the second. The value returned by hash-table-for-each is undefined.

Note: The order of application of proc is unspecified.

    (let ((h   (make-hash-table))
          (sum 0))
      (hash-table-put! h 'foo 2)
      (hash-table-put! h 'bar 3)
      (hash-table-for-each h (lambda (key value)
                               (set! sum (+ sum value))))
      sum)
              ⟹ 5

(hash-table-map hash proc)  procedure

Proc must be a procedure taking two arguments. Hash-table-map calls proc on each entry in hash, with the entry’s key as the first argument and the entry’s value as the second. The result of hash-table-map is a list of the values returned by proc, in unspecified order.

Note: The order of application of proc is unspecified.

    (let ((h (make-hash-table)))
      (dotimes (i 5)
        (hash-table-put! h i (number->string i)))
      (hash-table-map h (lambda (key value)
                           (cons key value))))
              ⟹ ((0 . "0") (3 . "3") (2 . "2") (1 . "1") (4 . "4"))

(hash-table->list hash)  procedure

hash-table->list returns an “association list” built from the entries in hash. Each entry in hash will be represented as a pair whose car is the entry’s key and whose cdr is its value.

Note: The order of pairs in the resulting list is unspecified.

    (let ((h (make-hash-table)))
      (dotimes (i 5)
        (hash-table-put! h i (number->string i)))
      (hash-table->list h))
              ⟹ ((0 . "0") (3 . "3") (2 . "2") (1 . "1") (4 . "4"))

(hash-table-stats hash)  procedure

Hash-table-stats returns a string with overall information about hash, such as the number of entries it contains, the number of buckets in its hash array, and the utilization of the buckets.

6.20 Regular expressions

Regular expressions are first class objects in STk. A regular expression is created by the string->regexp procedure. Matching a regular expression against a string is simply done by applying a previously created regular expression to this string. Regular expressions are implemented using code in the Henry Spencer’s package, and much of the description of regular expressions below is copied from his manual.

(string->regexp string)  procedure

String->regexp compiles the string and returns the corresponding regular expression. Matching a regular expression against a string is done by applying the result of string->regexp to this string. This application yields a list of integer couples if a matching occurs; it returns #f otherwise. Those integers correspond to indexes in the string which match the regular expression. A regular expression is zero or more branches, separated by “|”. It matches anything that matches one of the branches. A branch is zero or more pieces, concatenated. It matches a match for the first, followed by a match for the second, etc. A piece is an atom possibly followed by “*”, “+”, or “?”. An atom followed by “*” matches a sequence of 0 or more matches of the atom. An atom followed by “+” matches a sequence of 1 or more matches of the atom. An atom followed by “?” matches a match of the atom, or the null string.

An atom is a regular expression in parentheses (matching a match for the regular expression), a range (see below), “.” (matching any single character), “^” (matching the null string at the beginning of the input string), “$” (matching the null string at the end of the input string), a “\” followed by a single character (matching that character), or a single character with no other significance (matching that character). A range is a sequence of characters enclosed in “[]”. It normally matches any single character from the sequence. If the sequence begins with “^”, it matches any single character not from the rest of the sequence. If two characters in the sequence are separated by “-”, this is shorthand for the full list of ASCII characters between them (e.g. “[0-9]” matches any decimal digit). To include a literal “]” in the sequence, make it the first character (following a possible “^”). To include a literal “-”, make it the first or last character. In general there may be more than one way to match a regular expression to an input string. Considering only the rules given so far could lead to ambiguities. To resolve those ambiguities, the generated regular expression chooses among alternatives using the rule “first then longest”. In other words, it considers the possible matches in order working from left to right across the input string and the pattern, and it attempts to match longer pieces of the input string before shorter ones. More specifically, the following rules apply in decreasing order of priority:

1. If a regular expression could match two different parts of an input string then it will match the one that begins earliest.

2. If a regular expression contains “|” operators then the leftmost matching sub-expression is chosen.

3. In “*”, “+”, and “?” constructs, longer matches are chosen in preference to shorter ones.

4. In sequences of expression components the components are considered from left to right.

    (define r1 (string->regexp "abc"))
    (r1 "xyz")  ⟹ #f
    (r1 "12abc345")  ⟹ ((2 5))
    (define r2 (string->regexp "[a-z]+"))
    (r2 "12abc345")  ⟹ ((2 5))

If the regular expression contains parenthesis, and if there is a match, the result returned by the application will contain several couples of integers. First couple will be the indexes of the first longest substring which match the regular expression. Subsequent couples, will be the indexes of all the sub-parts of this regular expression, in sequence.

    (define r3 (string->regexp "(a*)(b*)c"))
    (r3 "abc")  ⟹ ((0 3) (0 1) (1 2))
    (r3 "c")  ⟹ ((0 1) (0 0) (0 0))
    ((string->regexp "([a-z]+),([a-z]+)") "XXabcd,eXX")
                                ⟹ ((2 8) (2 6) (7 8))

(regexp? obj )  procedure

Returns #t if obj is a regular expression created by string->regexp; otherwise returns #f.

    (regexp? (string->regexp "[a-zA-Z][a-zA-Z0-9]*"))
                                ⟹ #t

(regexp-replace pattern string substitution)  procedure
(regexp-replace-all pattern string substitution)  procedure

Regexp-replace matches the regular expression pattern against string. If there is a match, the portion of string which match pattern is replaced by the substitution string. If there is no match, regexp-replace returns string unmodified. Note that the given pattern could be here either a string or a regular expression. If pattern contains strings of the form “\n”, where n is a digit between 1 and 9, then it is replaced in the substitution with the portion of string that matched the n-th parenthesized subexpression of pattern. If n is equal to 0, then it is replaced in substitution with the portion of string that matched pattern.

    (regexp-replace "a*b" "aaabbcccc" "X")
                                ⟹ "Xbcccc"
    (regexp-replace (string->regexp "a*b") "aaabbcccc" "X")
                                ⟹ "Xbcccc"
    (regexp-replace "(a*)b" "aaabbcccc" "X\\1Y")
                                ⟹ "XaaaYbcccc"
    (regexp-replace "(a*)b" "aaabbcccc" "X\\0Y")
                                ⟹ "XaaabYbcccc"
    (regexp-replace "([a-z]*) ([a-z]*)" "john brown" "\\2 \\1")
                                ⟹ "brown john"

Regexp-replace replaces the first occurence of pattern in string. To replace all the occurences

of the pattern, use regexp-replace-all
    (regexp-replace "a*b" "aaabbcccc" "X")
                                ⟹ "Xbcccc"
    (regexp-replace-all "a*b" "aaabbcccc" "X")
                                ⟹ "XXcccc"

6.21 Pattern matching

Pattern matching is a key feature of most modern functional programming languages since it allows clean and secure code to be written. Internally, “pattern-matching forms” should be translated (compiled) into cascades of “elementary tests” where code is made as efficient as possible, avoiding redundant tests; the STk “pattern matching compiler” provides this^{7}. The technique used is described in details in [9], and the code generated can be considered optimal due to the way this “pattern compiler” was obtained. The “pattern language” allows the expression of a wide variety of patterns, including:

7The “pattern matching compiler” has been written by Jean-Marie Geffroy and is part of the Manuel Serrano’s Bigloo compiler[8] since several years. The code (and documentation) included in STk has been stolen from the Bigloo package v1.9 (the only difference between both package is the pattern matching of structures whisch is absent in STk).

• Non-linear patterns: pattern variables can appear more than once, allowing comparison of subparts of the datum (through eq?)

• Recursive patterns on lists: for example, checking that the datum is a list of zero or more as followed by zero or more bs.

• Pattern matching on lists as well as on vectors.

Pattern Matching Facilities

Only two special forms are provided for this: match-case and match-lambda and these also exist, for example, in Andrew Wright and Bruce Duba’s [10] pattern matching package.

(match-case ⟨key⟩ ⟨clause_{1}⟩ ⟨clause_{2}⟩…)  syntax

In this form, ⟨key⟩ may be any expression and each ⟨clause⟩ has the form

    (<pat> <expression_{1}> <expression_{2}> ...)

A match-case expression is evaluated as follows. <key> is evaluated and the result is compared with each successive patterns. If the pattern in some <clause> yields a match, then the expressions in that <clause> are evaluated from left to right in an environment where the pattern variables are bound to the corresponding subparts of the datum, and the result of the last expression in that <clause> is returned as the result of the match-case expression. If no <pat> in any <clause> matches the datum, then, if there is an else clause, its expressions are evaluated and the result of the last is the result of the whole match-case expression; otherwise the result of the match-case expression is unspecified. The equality predicate used is eq?.

    (match-case '(a b a)
      ((?x ?x) 'foo)
      ((?x ?- ?x) 'bar))
              ⟹ bar

(match-lambda ⟨clause_{1}⟩ ⟨clause_{2}⟩…)  syntax

The form match-lambdaexpands into a lambda-expression expecting an argument which, once applied to an expression, behaves exactly like a match-case expression.

    ((match-lambda
        ((?x ?x) 'foo)
        ((?x ?- ?x) 'bar)) 'bar)
              ⟹ bar

<pattern> ⟶  Matches:
<atom> ∥ (kwote <atom>) any expression eq? to <atom>
∥ (and <pat_{1} > … <pat_{n} >) if all of <pat_{i} > match
∥ (or <pat> … <pat_{n} >)  if any of <pat_{1} > through <pat_{n} >match
∥ (not <pat>)  if <pat> doesn’t match
∥ (? <predicate>)  if <predicate> is true
∥ (<pat_{1} > ...^{8} <pat_{n} >) a list of n elements
∥ <pat> \underline{...} ^{9}  a (possibly empty) repetition of <pat> in a list.
∥ #(<pat> ...<pat_{n} >)  a vector of n elements
∥ ?<identifier>  anything, and binds identifier as a variable
∥ ?-  anything
∥ ??-  any (possibly empty) repetition of anything in a list
∥ ???-  any end of list

Table 3: Pattern Syntax

The pattern language The syntax is presented in Table 3. It is described below in the same way (and nearly in the same words) as in [10].

Note: and, or, not, check and kwote must be quoted in order to be treated as literals. This is the only justification for having the kwote pattern since, by convention, any atom which is not a keyword is quoted.

Explanations through examples • ?- matches any s-expr

• a matches the atom 'a.

• ?a matches any expression, and binds the variable a to this expression.

• (? integer?) matches any integer

• (a (a b)) matches the only list '(a (a b)).

• ???- can only appear at the end of a list, and always succeeds. For instance, (a ???-) is equivalent to (a . ?-).

• when occurring in a list, ??- matches any sequence of anything: (a ??- b) matches any list whose car is a and last car is b.

• (a ...) matches any list of a’s, possibly empty.

• (?x ?x) matches any list of length 2 whose car is eq to its cadr

• ((and (not a) ?x) ?x) matches any list of length 2 whose car is not eq to 'a but is eq to its cadr

• #(?- ?- ???-) matches any vector whose length is at least 2.

Note: ??- and ... patterns can not appear inside a vector, where you should use ???-: For example, #(a ??- b) or #(a...) are invalid patterns, whereas #(a ???-) is valid and matches any vector whose first element is the atom a.

6.22 Processes

STk provides access to Unix processes as first class objects. Basically, a process contains four informations: the standard Unix process identification (aka PID) and the three standard files of the process.

(run-process command p_{1} p_{2} p_{3} …)  procedure

run-process creates a new process and run the executable specified in command. The p correspond to the command line arguments. The following values of p have a special meaning:

• :input permits to redirect the standard input file of the process. Redirection can come from a file or from a pipe. To redirect the standard input from a file, the name of this file must be specified after :input. Use the special keyword :pipe to redirect the standard input from a pipe.

• :output permits to redirect the standard output file of the process. Redirection can go to a file or to a pipe. To redirect the standard output to a file, the name of this file must be specified after :output. Use the special keyword :pipe to redirect the standard output to a pipe.

• :error permits to redirect the standard error file of the process. Redirection can go to a file or to a pipe. To redirect the standard error to a file, the name of this file must be specified after :error. Use the special keyword :pipe to redirect the standard error to a pipe.

• :wait must be followed by a boolean value. This value specifies if the process must be run asynchronously or not. By default, the process is run asynchronously (i.e. :wait is

     #f).

• :host must be followed by a string. This string represents the name of the machine on which the command must be executed. This option uses the external command rsh. The shell variable PATH must be correctly set for accessing it without specifying its abolute path.

The following example launches a process which execute the Unix command ls with the arguments -l and /bin. The lines printed by this command are stored in the file /tmp/X

    (run-process "ls" "-l" "/bin" :output "/tmp/X" :wait #f)

(process? process)  procedure

Returns #t if process is a process, otherwise returns #f.

(process-alive? process)  procedure

Returns #t if process if the process is currently running, otherwise returns #f.

(process-pid process)  procedure

Returns an integer value which represents the Unix identification (PID) of process.

(process-input process)  procedure

(process-output process)  procedure

(process-error process)  procedure

Returns the file port associated to the standard input, output or error of process, if it is redirected in (or to) a pipe; otherwise returns #f. Note that the returned port is opened for reading when calling process-output or process-error; it is opened for writing when

calling process-input.

(process-wait process)  procedure

Process-wait stops the current process until process completion. Process-wait returns #f when process is already terminated; it returns #t otherwise.

(process-exit-status process)  procedure

Process-exit-status returns the exit status of process if it has finished its execution; returns #f otherwise.

(process-send-signal process n)  procedure

Send the signal whose integer value is n to process. Value of n is system dependant. Use the defined signal constants to make your program indpendant of the running system (see 6.18). The result of process-send-signal is undefined.

(process-kill process)  procedure

Process-kill brutally kills process. The result of process-kill is undefined. This procedure is equivalent to

    (process-send-signal process |SIGTERM|)

(process-stop process)  procedure

(process-continue process)  procedure

Those procedures are only available on systems which support job control. Process-stop stops the execution of process and process-continue resumes its execution. They are equivalent to

    (process-send-signal process |SIGSTOP|)
    (process-send-signal process |SIGCONT|)

(process-list)  procedure

process-list returns the list of processes which are currently running (i.e. alive).

6.23 Sockets

STk defines sockets, on systems which support them, as first class objects. Sockets permits processes to communicate even if they are on different machines. Sockets are useful for creating client-server applications.

(make-client-socket hostname port-number)  procedure

make-client-socket returns a new socket object. This socket establishes a link between the running application listening on port port-number of hostname.

(socket? socket)  procedure

Returns #t if socket is a socket, otherwise returns #f.

(socket-host-name socket)  procedure

Returns a string which contains the name of the distant host attached to socket. If socket has been created with make-client-socket this procedure returns the official name of the distant machine used for connection. If socket has been created with make-server-socket, this function returns the official name of the client connected to the socket. If no client has used yet the socket, this function returns #f.

(socket-host-address socket)  procedure

Returns a string which contains the IP number of the distant host attached to socket. If socket has been created with make-client-socket this procedure returns the IP number of the distant machine used for connection. If socket has been created with make-server-socket, this function returns the address of the client connected to the socket. If no client has used yet the socket, this function returns #f.

(socket-local-address socket)  procedure

Returns a string which contains the IP number of the local host attached to socket.

(socket-port-number socket)  procedure

Returns the integer number of the port used for socket.

(socket-input socket)  procedure

(socket-output socket)  procedure

Returns the file port associated for reading or writing with the program connected with socket. If no connection has already been established, these functions return #f. The following example shows how to make a client socket. Here we create a socket on port 13 of the machine “kaolin.unice.fr”^{10}:

    (let ((s (make-client-socket "kaolin.unice.fr" 13)))
      (format #t  "Time is: ~A\n" (read-line (socket-input s)))
      (socket-shutdown  s))

10Port 13 is generally used for testing: making a connection to it permits to know the distant system’s idea of the time of day.

(make-server-socket)  procedure
(make-server-socket port-number)  procedure

make-server-socket returns a new socket object. If port-number is specified, the socket is listening on the specified port; otherwise, the communication port is chosen by the system.

(socket-accept-connection socket)  procedure

socket-accept-connection waits for a client connection on the given socket. If no client is already waiting for a connection, this procedure blocks its caller; otherwise, the first connection request on the queue of pending connections is connected to socket. This procedure must be called on a server socket created with make-server-socket. The result of

socket-accept-connection is undefined.

The following example is a simple server which waits for a connection on the port 1234^{11}. Once the connection with the distant program is established, we read a line on the input port associated to the socket and we write the length of this line on its output port.

    (let ((s (make-server-socket 1234)))
      (socket-accept-connection s)
      (let ((l (read-line (socket-input s))))
        (format (socket-output s) "Length is: ~A\n" (string-length l))
        (flush (socket-output s)))
      (socket-shutdown s))

(socket-shutdown socket)  procedure

(socket-shutdown socket close)  procedure

Socket-shutdown shutdowns the connection associated to socket. Close is a boolean; it indicates if the socket must be closed or not, when the connection is destroyed. Closing the socket forbids further connections on the same port with the socket-accept-connection procedure. Omitting a value for close implies the closing of socket. The result of socket-shutdown is undefined. The following example shows a simple server: when there is a new connection on the port number 1234, the server displays the first line sent to it by the client, discards the others and go back waiting for further client connections.

    (let ((s (make-server-socket 1234)))
      (let loop ()
        (socket-accept-connection s)
        (format #t "I've read: ~A\n" (read-line (socket-input s)))
        (socket-shutdown s #f)
        (loop)))

(socket-down? socket)  procedure

Returns #t if socket has been previously closed with socket-shutdown. It returns #f otherwise.

11Under Unix, you can simply connect to listening socket with the telnet command. With the given example, this can be achieved by typing the following command in a window shell:

$ telnet localhost 1234

(socket-dup socket)  procedure

Returns a copy of socket. The original and the copy socket can be used interchangeably. However, if a new connection is accepted on one socket, the characters exchanged on this socket are not visible on the other socket. Duplicating a socket is useful when a server must accept multiple simultaneous connections. The following example creates a server listening on port 1234. This server is duplicated and, once two clients are present, a message is sent on both connections.

    (define s1 (make-server-socket 1234))
    (define s2 (socket-dup s1))
    (socket-accept-connection s1)
    (socket-accept-connection s2)
    ;; blocks until two clients are present
    (display "Hello,\n" (socket-output s1))
    (display "world\n"  (socket-output s2))
    (flush (socket-output s1))
    (flush (socket-output s2))

(when-socket-ready socket handler)  procedure

(when-socket-ready socket)  procedure

Defines a handler for socket. The handler is a thunk which is executed when a connection is available on socket. If the special value #f is provided as handler, the current handler for socket is deleted. If a handler is provided, the value returned by when-socket-ready is undefined. Otherwise, it returns the handler currently associated to socket. This procedure, in conjunction with socket-dup permits to build multiple-clients servers which work asynchronously. Such a server is shown below.

    (define p (make-server-socket 1234))
    (when-socket-ready p
                      (let ((count 0))
                        (lambda ()
                          (set! count (+ count 1))
                          (register-connection (socket-dup p) count))))
    (define register-connection
      (let ((sockets '()))
        (lambda (s cnt)
          ;; Accept connection
          (socket-accept-connection s)
          ;; Save socket somewhere to avoid GC problems
          (set! sockets (cons s sockets))
          ;; Create a handler for reading inputs from this new connection
          (let ((in  (socket-input s))
                (out (socket-output s)))
            (when-port-readable in
                 (lambda ()
                   (let ((l (read-line in)))
                     (if (eof-object? l)
                         ;; delete current handler
                         (when-port-readable in #f)

Name Corresponding C type Corresponding Scheme

Type

:void            void                       None

:char

                 char

Scheme character or Scheme integer

:short

                 short int

Scheme integer

:ushort

                 unsigned short int  Scheme integer
:int

                 int

Scheme integer

:uint

                 unsigned integer  Scheme integer
:long

long integer Scheme integer

:ulong

                 unsigned long integer  Scheme integer

:float           float                      Scheme Real
:double          double                     Scheme Real

:static-ptr

pointer on a static area Scheme C-pointer object or

Scheme String

:dynamic-ptr or
(:void *)

pointer on a dynamic area

                 (mallocated)

Scheme C-pointer object or Scheme String

:string  or
(:char *)

char * (pointer on a dynamic

string)

Scheme C-pointer object or Scheme String

:boolean

                 int

Scheme boolean

Table 4: FFI predefined types

                    ;; Just write the line read on the socket
                    (begin
                      (format out "On #~A --> ~A\n" cnt l)
                      (flush out))))))))))

6.24 Foreign Function Interface

The STk Foreign Function Interface (FFI for short) has been defined to allow an easy access to functions written in C without needing to build C-wrappers and, consequently, without any need to write C code. Note that the FFI is very machine dependent and that it works only on a limited set of architectures^{12}. Moreover, since FFI allows very low level access, it is easy to crash the interpreter when using an external C function. The definition of an external function is done with the syntax define-external. This form takes as arguments a typed list of parameters and accepts several options to define the name of the function in the C world, the library which defines this function, … The type of the function result and the types of its arguments are defined in Table 4. This table lists the various keywords reserved for denoting types and their equivalence between the C and the Scheme worlds.

12In release 3.99, FFI is known to work on the following architectures : ix86 (but not yet MS Windows), Sun Sparc, HP 9000, SGI.

(define-external ⟨name⟩ ⟨parameters⟩ ⟨options⟩)  syntax

The form define-external binds a new procedure to ⟨name⟩. The arity of this new procedure is defined by the typed list of parameters given by ⟨parameters⟩. This parameters list is a list of couples whose first element is the name of the parameter, and the second one is is a keyword representing its type (see table for equivalence). All the types defined in Table 4, except :void, are allowed for the parameters of a foreign function. Define-external accepts several options:

• :return-type is used to define the type of the value returned by the foreign function. The type returned must be chosen in the types specified in the table. For instance:

        (define-external maximum((a :int) (b :int))
          :return-type :int)

defines the foreign function maximum which takes two C integers and returns an integer result. Omitting this option default to a result type equal to :void (i.e. the returned value is undefined).

• :entry-name is used to specify the name of the foreign function in the C world. If this option is omitted, the entry-name is supposed to be ⟨name⟩. For instance:

        (define-external minimum((a :int) (b :int))
          :return-type :int
          :entry-name  "min")

defines the Scheme function minimum whose application executes the C function called

min.

• :library-name is used to specify the library which contains the foreign-function. If necessary, the library is loaded before calling the C function. So,

        (define-external minimum((a :int) (b :int))
          :return-type  :int
          :entry-name   "min"
          :library-name "libminmax")

defines a function which will execute the function min located in the library libminmax.xx (where xx is the suffix used for shared libraries on the running system (generally so or

sl).

Hereafter, there are some commented definitions of external functions:

 (define-external isatty ((fd :int))
    :return-type :boolean)
 (define-external system ((cmd (:char *))) ;; or ((cmd :string))
    :return-type :int)
 (define-external malloc ((size :ulong))

    :return-type (void *))
 (define-external free ( (p (:void *) )) )

All these functions are defined in the C standard library, hence it is not necessary to specify

the :library-name option.

• istty is declared here as a function which takes an integer and returns a boolean (in fact, the value returned by the C function isatty is an int, but we ask here to the FFI system to translate this result as a boolean value in the Scheme world).

• system is a function which takes a string as parameter and returns an int. Note that the type of the parameter, can be specified as a (:char *) or :string, as indicated in Table 4.

• malloc is a function which takes one parameter (an unsigned long int and which returns a (:void *) (or :dynamic-ptr). Specifying that the result is a dynamic pointer (instead of a static one) means that we want that the Garbage Collector takes into account the area allocated by the C function malloc (i.e. if this area becomes no more accessible, the GC disposes it with the free function^{13}.

• free is a function which takes a dynamic pointer and deallocates the area it points. Since the definition of this function specifies no result type, it is supposed to be :void^{14}.

External functions can also have a variable number of parameters by using the standard Scheme dot notation. For instance,

(define-external printf ((format :string) . l)
 :return-type  :int)

defines a Scheme function with one or more parameters (the first one being a string). Of course, the parameters which constitute the variable parameters list must have a type which appears in the third column of Table 4. Some examples using the printf function:

 (printf "This is a %s test" "good")
         ⟹ displays   This is a good test
 (printf "char: '%c' Dec: '%04d' Hex '%04x'" #\space 100 100)
         ⟹ displays   char: ' ' Dec: '0100' Hex '0064'

Note: The types :dynamic-ptr, :static-ptr and :string are compatible when used for foreign function parameter. This gives a semantic which is similar to the one of C, where void * is a compatible with all other pointer types. However, differenciating those types is useful for converting the function return value to a proper Scheme type.

13Pointers defined with :dynamic-ptr are always unallocated with free. Consequently, areas allocated with another allocator than the standard one must be declared as :static-ptr and freed by hand

14Usage of malloc and free are for illustration purpose here. Their usage in a program must be avoided, if possible, because it can have interact badly with the way the interpreter manages memory or it can conduct to crashing programs if you don’t take care.

Note: When a function has a :return-type which is :string, :dynamic-ptr or :static-ptr, and

the return value is the C NULL pointer, the Scheme value returned by the function is, by convention, equal to #f. For instance, the GNU readline function allows line editing à la Emacs returns NULL when the user has typed an end of file. The following lines show how to make a simple shell-like toplevel using FFIs.

(define-external system ((var (:char *)))
:return-type :int)
(define-external readline ((prompt :string))
:library-name "libreadline"
:return-type  :string)
;; A Shell-like toplevel
(do ((l (readline "?> ") (readline "?> ")))
  ((not l))
(system l))

(external-exists? entry)  procedure

(external-exists? entry library)  procedure

Returns #t if entry is defined as an external symbol in library. If library is not provided the symbol is searched in the STk interpreter or in libraries that it uses. This function can be useful to define external functions conditionally:

(when (external-exists? "dup2")
   (define-external dup2 ((oldfd :int) (newfd :int))
      :return-type :int))

(c-string->string str)  procedure

STk strings are more general than C strings since they accept null character. c-string->string takes an area of characters built by a call to a foreign function (typically the result of a function returning a :static-ptr, :dynamic-ptr or :string) and convert it to a proper Scheme string.

(define-external sprintf ((str :string) (format :string) . l)
 :return-type  :int)
(let ((str (make-string 5 #\?)))
(sprintf str "%x" 100)
(cons str (C-string->string str)))
        ⟹ ("64\0??" . "64")

6.25 Miscellaneous

This section lists the primitives defined in STk that did not fit anywhere else.

(eval ⟨expr⟩)  syntax
(eval ⟨expr⟩ ⟨environment⟩)  syntax

Evaluates ⟨expr⟩ in the given environment. ⟨Environment⟩ may be omitted, in which case it defaults to the global environment.

    (define foo (let ((a 1)) (lambda () a)))
    (foo)  ⟹ 1
    (eval '(set! a 2) (procedure-environment foo))
    (foo)  ⟹ 2

(version)  procedure

returns a string identifying the current version of STk.

(machine-type)  procedure

returns a string identifying the kind of machine which is running the interpreter. The form

of the result is [os-name]-[os-version]-[processor-type].

(random n)  procedure

returns an integer in the range 0, n − 1 inclusive.

(set-random-seed! seed)  procedure

Set the random seed to the specified seed. Seed must be an integer which fits in a C long

int.

(eval-string string environment)  procedure

Evaluates the contents of the given string in the given environment and returns its result. If environment is omitted it defaults to the global environment. If evaluation leads to an error, the result of eval-string is undefined.

    (define x 1)
    (eval-string "(+ x 1)")
              ⟹ 2
    (eval-string "x" (let ((x 2)) (the-environment)))
              ⟹ 2

(read-from-string ⟨string⟩)  procedure

Performs a read from the given string. If string is the empty string, an end of file object is returned. If an error occurs during string reading, the result of read-from-string is undefined.

    (read-from-string "123 456")
              ⟹ 123
    (read-from-string "")
              ⟹ an eof object

(dump string)  procedure

Dump grabs the current continuation and creates an image of the current STk interpreter in the file whose name is string^{15}. This image can be used later to restart the interpreter from the saved state. See the STk man page about the -image option for more details.

Note: Image creation cannot be done if Tk is initialized.

(trace-var symbol thunk)  procedure

Trace-var call the given thunk when the value of the variable denoted by symbol is changed.

    (define x 1)
    (define y 0)
    (trace-var 'x (lambda () (set! y 1)))
    (set! x 2)
    (cons x y)
              ⟹ (2 . 1)

Note: Several traces can be associated with a single symbol. They are executed in reverse order to their definition. For instance, the execution of

    (begin
      (trace-var 'z (lambda () (display "One")))
      (trace-var 'z (lambda () (display "Two")))
      (set! z 10))

will display the string "Two" before the string "One" on the current output port.

(untrace-var symbol)  procedure

Deletes all the traces associated to the variable denoted by symbol.

(error string string_{1} obj_{2} …)  procedure

error prints the obj s according to the specification given in string on the current error port (or in an error window if Tk is initialized). The specification string follows the “tilde conventions” of format(see 6.10). Once the message is printed, execution returns to toplevel.

(gc)  procedure

Runs the garbage collector. See 6.18 for the signals associated to garbage collection.

15Image creation is not yet implemented on all systems. The current version (3.99) allows image dumping only on some platforms: SunOs 4.1.x, Linux 1, FreeBsd

(gc-stats)  procedure

Provides some statistics about current memory usage. This procedure is primarily for debugging the STk interpreter, hence its weird printing format.

(expand-heap n)  procedure

Expand the heap so that it will contains at least n cells. Normally, the heap automatically grows when more memory is needed. However, using only automatic heap growing is sometimes very penalizing. This is particularly true for programs which uses a lot of temporary data (which are not pointed by any a variable) and a small amount of global data. In this case, the garbage collector will be often called and the heap will not be automatically expanded (since most of the consumed heap will be reclaimed by the GC). This could be annoying specially for program where response time is critical. Using expand-heap permits to enlarge the heap size (which is set to 20000 cells by default), to avoid those continual calls to the GC.

(get-internal-info)  procedure

Returns a 7-length vector which contains the following informations:

0 total cpu used in milli-seconds

1 number of cells currently in use.

2 total number of allocated cells

3 number of cells used since the last call to get-internal-info

4 number of gc runs

5 total time used in the gc

6 a boolean indicating if Tk is initialized

(sort obj predicate)  procedure

Obj must be a list or a vector. Sort returns a copy of obj sorted according to predicate. Predicate must be a procedure which takes two arguments and returns a true value if the first argument is strictly “before” the second.

    (sort '(1 2 -4 12 9 -1 2 3) <)
                                ⟹ (-4 -1 1 2 2 3 9 12)
    (sort #("one" "two" "three" "four")
          (lambda (x y) (> (string-length x) (string-length y))))
                                ⟹ #("three" "four" "one" "two")

(uncode form)  procedure

When STk evaluates an expression it encodes it so that further evaluations of this expression will be more efficient. Since encoded forms are generally difficult to read, uncode can be used to (re-)obtain the original form.

    (define (foo a b)
       (let ((x a) (y (+ b 1))) (cons x y)))
    (procedure-body foo)
                                ⟹ (lambda (a b)
                           (let ((x a) (y (+ b 1))) (cons x y)))
    (foo 1 2)  ⟹ (1 . 3)
    (procedure-body foo)
                                ⟹ (lambda (a b)
                           (#let (x y)
                                 (#<local a @0,0)>
                                  (#<global +> #<local b @0,1)> 1))
                             (#<global cons> #<local x @0,0)>
                                             #<local y @0,1)>)))
    (uncode (procedure-body foo))
                                ⟹ (lambda (a b)
                            (let ((x a) (y (+ b 1))) (cons x y)))

Note: When a macro has been directly expanded into the macro call code, it is not possible to retrieve the original macro call. Set *debug* to #t to avoid macro expansion in-lining.

(time ⟨expr⟩)  macro

Evaluates the expression ⟨expr⟩ in the current environment. Prints the elapsed CPU time and the number of conses used before returning the result of this evaluation.

(apropos symbol)  procedure

Apropos returns a list of symbol whose print name contains the characters of symbol. Symbols are searched for in the current environment.

    (apropos 'cadd)
              ⟹ (caddar caddr cadddr)

(inspect obj )  procedure

Inspect permits to graphically inspect an object. The first call of this procedure creates a top level window containing the object to inspect and its current value. If the inspector window is already on screen, obj will be appended to the list of inspected objects. The inspector window contains menus which permit to call the viewer or detailer on each inspected object. See the on-line documentation for further details. A view of the general inspector is given in figure 1.

Note: Tk must be initialized to use inspect.

(view obj )  procedure

View permits to obtain a graphical representation of an STk object. The type of representation depends on the type of the viewed object. Here again, menus are provided to switch to

62  STk Reference Manual

Figure 1: A view of the Inspector Figure 2: A view of the Viewer

Figure 3: A view of the Detailer

the inspector or to the detailer. See the on-line documentation for more details. A snapshot of the viewer is given in figure 2.

Note: Tk must be initialized to use view.

(detail obj )  procedure

detail permits to display the fields of a composite Scheme object. The type of detailer depends on the type of the composite object detailed. Here again, menus are provided to go to the inspector or to the viewer. See the on-line documentation for more details. Figure 3 shows the detailer examining a tk-command.

Note: Tk must be initialized to use detail.

(quit retcode)  procedure
(quit)  procedure
(exit retcode)  procedure
(exit)  procedure

(bye retcode)  procedure
(bye)  procedure

Exits the STk interpreter with the specified integer return code. If omitted, the interpreter terminates with a return code of 0.

Part II Annexes

65

Appendix A

Using the Tk toolkit

When STk detects that a tk-command must be called, parameters are processed to be recognized by the corresponding toolkit function. Since the Tk toolkit is left (mostly) unmodified, all its primitives “think” there is a running Tcl interpreter behind the scene. Consequently, to work with the Tk toolkit, a little set of rewriting rules must be known. These rules are described hereafter.

Note: This appendix is placed here to permit an STk user to make programs with the original Tcl/Tk documentation by hand. In no case will it substitute to the abundant Tcl/Tk manual pages nor to the excellent book by J. Ousterhout[11]

1 Calling a Tk-command

Since Tcl uses strings to communicate with the Tk toolkit, parameters to a Tk-command must be translated to strings before calling the C function which implement it. The following conversions are done, depending on the type of the parameter that STk must give to the toolkit:

symbol: the print name of the symbol;

number: the external representation of the number expressed in radix 10;

string: no conversion;

keyword: the print name of the keyword where the initial semicolon has been replaced by a dash (“-”);

boolean: the string ”0” if #f and ”1” if #t

tk-command: the name of the tk-command

closure: the address of the closure using the representation shown in 6.17.

otherwise: the external “slashified” version of the object.

As an example, let us make a button with a label containing the string "Hello, word". According the original Tk/Tcl documentation, this can be done in Tcl with

    button .hello -text "Hello, world"

Following the rewriting rules expressed above, this can be done in STk with

    (button '.hello '-text "Hello, world")

This call defines a new widget object which is stored in the STk variable .hello. This object can be used as a procedure to customize our button. For instance, setting the border of this button to 5 pixels wide and its background to gray would be done in Tcl with

    .hello configure -border 5 -background gray

In STk this would be expressed as

    (.hello 'configure '-border 5 '-background "gray")

Since keyword colon is replaced by a dash when a Tk-command is called, this expression could also have been written as:

    (.hello 'configure :border 5 :background "gray")

2 Associating Callbacks to Tk-commands

Starting with version 3.0, STk callbacks are Scheme closures^{1}. Apart scroll commands, callbacks are Schemes procedures without parameter. Suppose for example, that we want to associate a command with the previous .hello button. In Tcl, such a command can be expressed as

    .hello configure -command {puts stdout "Hello, world"; destroy .}

In STk, we can write

    (.hello 'configure :command (lambda ()
                                       (display "Hello, world\n")
                                       (destroy *root*)))

When the user will press the mouse left button, the closure associated to the :command option will be evaluated in the global environment. Evaluation of the given closure will display the message and call the destroy Tk-command.

Note: The root widget is denoted “.” in Tcl. This convention is ambiguous with the dotted pair convention and the dot must be quoted to avoid problems. Since this problem arises so often, the variable *root* has been introduced in STk to denote the Tk main window.

Managing Widget Scrollbars

When using scrollbars, Tk library passes parameters to the widget associated to the scrollbar (and vice versa). Let us look at a text widget with an associated scrollbar. When the scrollbar is moved, the command of the associated widget is invoked to change its view. On the other side, when browsing the content of the text widget (with arrows for example), the scrollbar is updated by calling it’s associated closure. Tk library passes position informations to scrolling closures. This informations are the parameters of the closure. Hereafter is an example implementing a text widget with a scrollbar (see the help pages for details):

1Old syntax for callbacks (i.e. strings) is always supported but its use is deprecated.

    (text '.txt :yscrollcommand  (lambda l (apply .scroll 'set l)))
    (scrollbar '.scroll :command (lambda l (apply .txt 'yview l)))
    (pack .txt :side "left")
    (pack .scroll :fill "y" :expand #t :side "left")

3 Tk bindings

Bindings are Scheme closures

The Tk bind command associates Scheme scripts with X events. Starting with version 3.0 those scripts must be Scheme closures^{2}. Binding closures can have parameters. Those parameters are one char symbols (with the same conventions than the Tcl % char, see the bind help page for details). For instance, the following Tcl script

    bind .w <ButtonPress-3> {puts "Press on widget %W at position %x %y"}

can be translated into

    (bind .w "<ButtonPress-3>"
             (lambda (|W| x y)
                (format #t "Press on widget ~A at position ~A ~A\n" |W| x y)))

Note: Usage of verticals bars for the W symbol is necessary here because the Tk toolkit is case sensitive (e.g. W in bindings is the path name of the window to which the event was reported, whereas w is the width field from the event.

Bindings are chained

In Tk4.0 and later, bindings are chained since it is possible for several bindings to match a given X event. If the bindings are associated with different tags, then each of the bindings will be executed, in order. By default, a class binding will be executed first, followed by a binding for the widget, a binding for its toplevel, and an all binding. The bindtags command may be used to change this order for a particular window or to associate additional binding tags with the window (see corresponding help page for details). If the result of closure in the bindings chain is the symbol break, the next closures of the chain are not executed. The example below illustrates this:

    (pack (entry '.e))
    (bind .e "<KeyPress>" (lambda (|A|)
                            (unless (string->number |A|) 'break)))

Bindings for the entry .e are executed before those for its class (i.e. Entry). This allows us to filter the characters which are effectively passed to the .e widget. The test in this binding closure breaks the chain of bindings if the typed character is not a digit. Otherwise, the following binding, the one for the Entry class, is executed and inserts the character typed (a digit). Consequently, the simple previous binding makes .e a controlled entry which only accepts integer numbers.

2Old syntax for bindings (i.e. strings) is no more supported. Old bindings scripts must hence be rewritten.

70  STk Reference Manual

Appendix B

Differences with R4RS

This appendix summarizes the main differences between the STk Scheme implementation and the language described in R^{4}RS.

1 Symbols

STk symbol syntax has been augmented to allow case significant symbols. This extension is discussed in 6.4. STk also defines some symbols in the global environment which are described below:

• *debug*. Setting *debug* to #t prevents macro inlining and expression recoding (see 6.25).

• *gc-verbose*. If *gc-verbose* is #t, a message will be printed before and after each run of garbage collector. The message is printed on the standard error stream.

• *load-verbose*. If *load-verbose* is #t, the absolute path name of each loaded file is printed before its effective reading. File names are printed on the standard error stream.

• *load-path* must contain a list of strings. Each string is taken as a directory path name in which a file will be searched for loading. This variable can be set automatically from the STK LOAD PATH shell variable. See stk(1) for more details.

• *load-suffixes* must contain a list of strings. When the system try to load a file in a given directory (according to *load-path* value), it will first try to load it without suffix. If this file does not exist, the system will sequentially try to find the file by appending each suffix of this list. A typical value for this variable may be ("stk"

"stklos" "scm" "so").

• *argc* contains the number of arguments (0 if none), not including interpreter options. See stk(1) for more details.

• *argv* contains a Scheme list whose elements are the arguments (not including the interpreter options), in order, or an empty list if there are no arguments. See stk(1) for more details.

• *program-name* contains the file name specified with the -file option, if present. Otherwise, it contains the name through which the interpreter was invoked. See stk(1) for more details.

• *print-banner*. If *print-banner* is #f, the usual copyright message is not displayed when the interpreter is started.

• *stk-library* contains the path name of the installation directory of the STk library. This variable can be set automatically from the STK LIBRARY shell variable. See stk(1) for more details.

The following symbols are defined only when Tk is loaded:

• *root* designates the Tk main window (see A-2). This variable is not set if the Tk toolkit is not initialized.

• *help-path* must contain a list of strings. Each string is taken as a directory path name in which documentation files are searched. This variable can be set automatically from the STK HELP PATH shell variable. See stk(1) for more details.

• *image-path* must contain a list of strings. Each string is taken as a directory path name in which images are searched by the function make-image. This variable can be set automatically from the STK IMAGE PATH shell variable. See stk(1) and make-image(n) for more details.

• *root* designates the Tk main window (see A-2). This variable is not set if the Tk toolkit is not initialized.

• *start-withdrawn*. If *start-withdrawn* is not false, the *root* window is not mapped on screen until its first sub-window is packed or some action is asked to the window manager for it.

• *tk-version* is a string which contains the version number of the Tk toolkit used by STk.

• *tk-patch-level* is a string which contains the version and patch level of the Tk toolkit used by STk.

Furthermore, STk also defines the following procedures in the global environment:

• report-error. This procedure is called by the error system to display the message error. This procedure is described in report-error(n)

• repl-display-prompt. This procedure is called when the system is run interactively before reading a sexpr to evaluate to display a prompt. This procedure is described in

repl-display-prompt(n).

• repl-display-result. This procedure is called when the system is run interactively after the evaluation of a sexpr to write the result. This procedure is described in

repl-display-result(n).

2 Types

STk implements all the types defined as mandatory in R^{4}RS. However, complex numbers and rational numbers (which are defined but not required in R^{4}RS) are not implemented. The lack of these types implies that some functions of R^{4}RS are not defined. Some types which are not defined in R^{4}RS are implemented in STk. Those types are listed below:

• input string port type (6.10)

• output string port type (6.10)

• keyword type (6.11)

• Tk command type (6.12)

• environment type (6.14)

• macro type (6.15)

• address type (6.17)

• hash table type (6.19)

• Regular expression type (6.20)

• process type (6.22)

• socket type (6.23)

3 Procedures

The following procedures are required by R^{4}RS and are not implemented in the STk interpreter.

• transcript-off

• transcript-on Transcript-off and transcript-on can be simulated with various Unix tools such as script

or fep.

The following procedures are not implemented in the STk interpreter whereas they are defined in R^{4}RS (but not required). They are all related to complex or rational numbers.

• numerator

• denominator

• rationalize

• make-rectangular

• make-polar • real-part

• imag-part

• magnitude

• angle

Appendix C

An introduction to STklos

1 Introduction

STklos is the object oriented layer of STk. Its implementation is derived from version 1.3 of the Gregor Kickzales Tiny Clos package [12]. However, it has been extended to be as close as possible to CLOS, the Common Lisp Object System[6]. Some features of STklos are also issued from Dylan[13] or SOS[14]. Briefly stated, the STklos extension gives the user a full object oriented system with metaclasses, multiple inheritance, generic functions and multi-methods. Furthermore, the whole implementation relies on a true meta object protocol, in the spirit of the one defined for CLOS[15]. This model has also been used to embody the predefined Tk widgets in a hierarchy of STklos classes. This set of classes permits to simplify the core Tk usage by providing homogeneous accesses to widget options and by hiding the low level details of Tk widgets, such as naming conventions. Furthermore, as expected, using of objects facilitates code reuse and definition of new widgets classes. The purpose of this appendix is to introduce briefly the STklos package and in no case will it replace the STklos reference manual (which needs to be urgently written now …). In particular, methods relative to the meta object protocol and access to the Tk toolkit will not be described here.

2 Class definition and instantiation

2.1 Class definition

A new class is defined with the define-class macro. The syntax of define-class is close

to CLOS defclass:

(define-class class (⟨superclass_{1}⟩ ⟨superclass_{2}⟩...) (⟨slot description_{1}⟩ ⟨slot description_{2}⟩...) ⟨metaclass option⟩)

The ⟨metaclass option⟩ will not be discussed in this appendix. The ⟨superclass⟩es list specifies the super classes of class (see 3 for more details). A ⟨slot description⟩ gives the name of a slot and, eventually, some “properties” of this slot (such as its initial value, the function which permit to access its value, …). Slot descriptions will be discussed in 3.3.

As an exemple, consider now that we have to define a complex number. This can be done with the following class definition:

    (define-class  <complex> (<number>)
       (r i))

This binds the symbol <complex> to a new class whose instances contain two slots. These slots are called r an i and we suppose here that they contain respectively the real part and the imaginary part of a complex number. Note that this class inherits from <number> which is a pre-defined class (<number> is the super class of the <real> and <integer> pre-defined classes).^{1}.

3 Inheritance

3.1 Class hierarchy and inheritance of slots

Inheritance is specified upon class definition. As said in the introduction, STklos supports multiple inheritance. Hereafter are some classes definition:

    (define-class A () (a))
    (define-class B () (b))
    (define-class C () (c))
    (define-class D (A B) (d a))
    (define-class E (A C) (e c))
    (define-class F (D E) (f))

A, B, C have a null list of super classes. In this case, the system will replace it by the list which only contains <object>, the root of all the classes defined by define-class. D, E, F use multiple inheritance: each class inherits from two previously defined classes. Those class definitions define a hierarchy which is shown in Figure 1. In this figure, the class <top> is also shown; this class is the super class of all Scheme objects. In particular, <top> is the super class of all standard Scheme types. The set of slots of a given class is calculated by “unioning” the slots of all its super class. For instance, each instance of the class D, defined before will have three slots (a, b and d). The slots of a class can be obtained by the class-slots primitive. For instance,

    (class-slots A)
              ⟹ (a)
    (class-slots E)
              ⟹ (a e c)
    (class-slots F)
              ⟹ (d a b c f)

Note: The order of slots is not significant.

1With this definition, a <real> is not a <complex> since <real> inherits from <number> rather than <complex>. In practice, inheritance could be modified a posteriori, if needed. However, this necessitates some knowledge of the meta object protocol and it will not be shown in this document

Figure C.1: A class hierarchy

3.2 Instance creation and slot access

Creation of an instance of a previously defined class can be done with the make procedure. This procedure takes one mandatory parameter which is the class of the instance which must be created and a list of optional arguments. Optional arguments are generally used to initialize some slots of the newly created instance. For instance, the following form

    (define c (make <complex>))

will create a new <complex> object and will bind it to the c Scheme variable. Accessing the slots of the new complex number can be done with the slot-ref and the slot-set! primitives. Slot-set! primitive permits to set the value of an object slot and slot-ref permits to get its value.

    (slot-set! c 'r 10)
    (slot-set! c 'i 3)
    (slot-ref c 'r)
              ⟹ 10
    (slot-ref c 'i)
              ⟹ 3

Using the describe generic function is a simple way to see all the slots of an object at one time: this function prints all the slots of an object on the standard output. For instance, the expression

    (describe c)

will print the following informations on the standard output:

#[<complex> 122398] is an instance of class <complex>
Slots are:
    r = 10
    i = 3

3.3 Slot description

When specifying a slot, a set of options can be given to the system. Each option is specified with a keyword. The list of authorised keywords is given below:

• :initform permits to supply a default value for the slot. This default value is obtained by evaluating the form given after the :initform in the global environment.

• :init-keyword permits to specify the keyword for initializing a slot. The init-keyword may be provided during instance creation (i.e. in the make optional parameter list). Specifying such a keyword during instance initialization will supersede the default slot initialization possibly given with :initform.

• :getter permits to supply the name for the slot getter. The name binding is done in the global environment.

• :setter permits to supply the name for the slot setter. The name binding is done in the global environment.

• :accessor permits to supply the name for the slot accessor. The name binding is done in the global environment. An accessor permits to get and set the value of a slot. Setting the value of a slot is done with the extended version of set!.

• :allocation permits to specify how storage for the slot is allocated. Three kinds of allocation are provided. They are described below:

– :instance indicates that each instance gets its own storage for the slot. This is the default.

– :class indicates that there is one storage location used by all the direct and indirect instances of the class. This permits to define a kind of global variable which can be accessed only by (in)direct instances of the class which defines this slot.

– :virtual indicates that no storage will be allocated for this slot. It is up to the user to define a getter and a setter function for this slot. Those functions must be defined with the :slot-ref and :slot-set! options. See the example below.

To illustrate slot description, we shall redefine the <complex> class seen before. A definition could be:

(define-class <complex> (<number>)
  ((r :initform 0 :getter get-r :setter set-r! :init-keyword :r)
   (i :initform 0 :getter get-i :setter set-i! :init-keyword :i)))

With this definition, the r and i slot are set to 0 by default. Value of a slot can also be specified by calling make with the :r and :i keywords. Furthermore, the generic functions get-r and set-r! (resp. get-i and set-i!) are automatically defined by the system to read and write the r (resp. i) slot.

    (define c1 (make <complex> :r 1 :i 2))
    (get-r c1)
              ⟹ 1
    (set-r! c1 12)
    (get-r c1)
              ⟹ 12
    (define c2 (make <complex> :r 2))
    (get-r c2)
              ⟹ 2
    (get-i c2)
              ⟹ 0

Accessors provide an uniform access for reading and writing an object slot. Writing a slot is done with an extended form of set! which is close to the Common Lisp setf macro. So, another definition of the previous <complex> class, using the :accessor option, could be:

    (define-class <complex> (<number>)
       ((r :initform 0 :accessor real-part :init-keyword :r)
        (i :initform 0 :accessor imag-part :init-keyword :i)))

Using this class definition, reading the real part of the c complex can be done with:

    (real-part c)

and setting it to the value contained in the new-value variable can be done using the extended form of set!.

    (set! (real-part c) new-value)

Suppose now that we have to manipulate complex numbers with rectangular coordinates as well as with polar coordinates. One solution could be to have a definition of complex numbers which uses one particular representation and some conversion functions to pass from one representation to the other. A better solution uses virtual slots. A complete definition of the <complex> class using virtual slots is given in Figure 2.

This class definition implements two real slots (r and i). Values of the m and a virtual slots are calculated from real slot values. Reading a virtual slot leads to the application of the function defined in the :slot-ref option. Writing such a slot leads to the application of the function defined in the :slot-set! option. For instance, the following expression

    (slot-set! c 'a 3)

permits to set the angle of the c complex number. This expression conducts, in fact, to the evaluation of the following expression

(define-class <complex> (<number>)
(;; True slots use rectangular coordinates
(r :initform 0 :accessor real-part :init-keyword :r)
(i :initform 0 :accessor imag-part :init-keyword :i)
;; Virtual slots access do the conversion
(m :accessor magnitude :init-keyword :magn
:allocation :virtual
:slot-ref (lambda (o)
            (let ((r (slot-ref o 'r)) (i (slot-ref o 'i)))
              (sqrt (+ (* r r) (* i i)))))
:slot-set! (lambda (o m)
              (let ((a (slot-ref o 'a)))
                (slot-set! o 'r (* m (cos a)))
                (slot-set! o 'i (* m (sin a))))))
(a :accessor angle :init-keyword :angle
:allocation :virtual
:slot-ref (lambda (o)
            (atan (slot-ref o 'i) (slot-ref o 'r)))
:slot-set! (lambda(o a)
             (let ((m (slot-ref o 'm)))
                (slot-set! o 'r (* m (cos a)))
                (slot-set! o 'i (* m (sin a))))))))

Figure C.2: A <complex> number class definition using virtual slots

((lambda o m)
(let ((m (slot-ref o 'm)))
(slot-set! o 'r (* m (cos a)))
(slot-set! o 'i (* m (sin a))))
c 3)

A more complete example is given below:

(define c (make <complex> :r 12 :i 20))
(real-part c)
⟹ 12
(angle c)
⟹ 1.03037682652431
(slot-set! c 'i 10)
(set! (real-part c) 1)
(describe c)

⟹

#[<complex> 128bf8] is an instance of class <complex>
Slots are:
   r = 1
   i = 10
   m = 10.0498756211209
   a = 1.47112767430373

Since initialization keywords have been defined for the four slots, we can now define the make-rectangular and make-polar standard Scheme primitives.

    (define make-rectangular
       (lambda (x y) (make <complex> :r x :i y)))
    (define make-polar
       (lambda (x y) (make <complex> :magn x :angle y)))

3.4 Class precedence list

A class may have more than one superclass.^{2} With single inheritance (one superclass), it is easy to order the super classes from most to least specific. This is the rule:

Rule 1: Each class is more specific than its superclasses.

With multiple inheritance, ordering is harder. Suppose we have

    (define-class X ()
       ((x :initform 1)))
    (define-class Y ()
       ((x :initform 2)))
    (define-class Z (X Y)
       (...))

In this case, the Z class is more specific than the X or Y class for instances of Z. However, the :initform specified in X and Y leads to a problem: which one overrides the other? The rule in STklos, as in CLOS, is that the superclasses listed earlier are more specific than those listed later. So:

Rule 2: For a given class, superclasses listed earlier are more specific than those listed later.

These rules are used to compute a linear order for a class and all its superclasses, from most specific to least specific. This order is called the “class precedence list” of the class. Given those two rules, we can claim that the initial form for the x slot of previous example is 2 since the class X is placed before Y in class precedence list of Z. This two rules are not always enough to determine a unique order, however, but they give an idea of how things work. STklos algorithm for calculating the precedence list is a little simpler than the CLOS one described in [15] for breaking ties. Consequently the calculated class precedence list could be different. Taking the F class shown in Figure 1, the STklos calculated class precedence list is

     (f d e a b c <object> <top>)

whereas it would be the following list with a CLOS-like algorithm:

     (f d e a c b <object> <top>)

2This section is an adaptation of Jeff Dalton’s (J.Dalton@ed.ac.uk) Brief introduction to CLOS)

However, it is usually considered a bad idea for programmers to rely on exactly what the order is. If the order for some superclasses is important, it can be expressed directly in the class definition. The precedence list of a class can be obtained by the function class-precedence-list. This function returns a ordered list whose first element is the most specific class. For instance,

    (class-precedence-list B)
              ⟹ (#[<class> 12a248] #[<class> 1074e8] #[<class> 107498])

However, this result is not too much readable; using the function class-name yields a clearer result:

    (map class-name (class-precedence-list B))
              ⟹ (b <object> <top>)

4 Generic functions

4.1 Generic functions and methods

Neither STklos nor CLOS use the message mechanism for methods as most Object Oriented language do. Instead, they use the notion of generic function. A generic function can be seen as a methods “tanker”. When the evaluator requestd the application of a generic function, all the methods of this generic function will be grabbed and the most specific among them will be applied. We say that a method M is more specific than a method M’ if the class of its parameters are more specific than the M’ ones. To be more precise, when a generic funtion must be “called” the system will

1. search among all the generic function those which are applicable

2. sort the list of applicable methods in the “most specific” order

3. call the most specific method of this list (i.e. the first method of the sorted methods list).

The definition of a generic function is done with the define-generic macro. Definition of a new method is done with the define-method macro. Note that define-method automatically defines the generic function if it has not been defined before. Consequently, most of the time, the define-generic needs not be used. Consider the following definitions:

    (define-generic M)
    (define-method M((a <integer>) b) 'integer)
    (define-method M((a <real>) b) 'real)
    (define-method M(a b) 'top)

The define-generic call defines M as a generic function. Note that the signature of the generic function is not given upon definition, contrarily to CLOS. This will permit methods with different signatures for a given generic function, as we shall see later. The three next lines define methods for the M generic function. Each method uses a sequence of parameter specializers that specify when the given method is applicable. A specializer permits to indicate the class a parameter must belong to (directly or indirectly) to be applicable. If no speciliazer is given, the system defaults it to <top>. Thus, the first method definition is equivalent to

    (define-method M((a <integer>) (b <top>)) 'integer)

Now, let us look at some possible calls to generic function M:

    (M 2 3)
              ⟹ integer
    (M 2 #t)
              ⟹ integer
    (M 1.2 'a)
              ⟹ real
    (M #3 'a)
              ⟹ real
    (M #t #f)
              ⟹ top
    (M 1 2 3)
              ⟹ error (since no method exists for 3 parameters)

The preceding methods use only one specializer per parameter list. Of course, each parameter can use a specializer. In this case, the parameter list is scanned from left to right to determine the applicability of a method. Suppose we declare now

    (define-method M ((a <integer>) (b <number>)) 'integer-number)
    (define-method M ((a <integer>) (b <real>))   'integer-real)
    (define-method M (a (b <number>)) 'top-number)

In this case,

    (M 1 2)
              ⟹ integer-integer
    (M 1 1.0)
              ⟹ integer-real
    (M 1 #t)
              ⟹ integer
    (M 'a 1)
              ⟹ 'top-number

4.2 Next-method

When a generic function is called, the list of applicable methods is built. As mentioned before, the most specific method of this list is applied (see 4.1). This method may call the next method in the list of applicable methods. This is done by using the special form next-method. Consider the following definitions

    (define-method Test((a <integer>))  (cons 'integer (next-method)))
    (define-method Test((a <number>))   (cons 'number  (next-method)))
    (define-method Test(a)     (list 'top))

With those definitions,

    (Test 1)
              ⟹ (integer number top)
    (Test 1.0)
              ⟹ (number top)
    (Test #t)
              ⟹ (top)

4.3 Example

In this section we shall continue to define operations on the <complex> class defined in Figure 2. Suppose that we want to use it to implement complex numbers completely. For instance a definition for the addition of two complexes could be

    (define-method new-+ ((a <complex>) (b <complex>))
      (make-rectangular (+ (real-part a) (real-part b))
                        (+ (imag-part a) (imag-part b))))

To be sure that the + used in the method new-+ is the standard addition we can do:

    (define-generic new-+)
    (let ((+ +))
      (define-method new-+ ((a <complex>) (b <complex>))
        (make-rectangular (+ (real-part a) (real-part b))
                          (+ (imag-part a) (imag-part b)))))

The define-generic ensures here that new-+ will be defined in the global environment. Once this is done, we can add methods to the generic function new-+ which make a closure on the + symbol. A complete writing of the new-+ methods is shown in Figure 3.

We use here the fact that generic function are not obliged to have the same number of parameters, contrarily to CLOS. The four first methods implement the dyadic addition. The fifth method says that the addition of a single element is this element itself. The sixth method says that using the addition with no parameter always return 0. The last method takes an arbitrary number of parameters^{3}. This method acts as a kind of reduce: it calls the dyadic addition on the car of the list and on the result of applying it on its rest. To finish, the set! permits to redefine the + symbol to our extended addition.

To terminate our implementation (integration?) of complex numbers, we can redefine standard Scheme predicates in the following manner:

    (define-method complex? (c <complex>) #t)
    (define-method complex? (c)           #f)
    (define-method number? (n <number>) #t)
    (define-method number? (n)          #f)

3The third parameter of a define-method is a parameter list which follow the conventions used for lambda expressions. In particular it can use the dot notation or a symbol to denote an arbitrary number of parameters

STk Reference Manual  85

(define-generic new-+)
(let ((+ +))
(define-method new-+ ((a <real>) (b <real>)) (+ a b))
(define-method new-+ ((a <real>) (b <complex>))
  (make-rectangular (+ a (real-part b)) (imag-part b)))
(define-method new-+ ((a <complex>) (b <real>))
  (make-rectangular (+ (real-part a) b) (imag-part a)))
(define-method new-+ ((a <complex>) (b <complex>))
  (make-rectangular (+ (real-part a) (real-part b))
                    (+ (imag-part a) (imag-part b))))
(define-method new-+ ((a <number>))  a)
(define-method new-+ () 0)
(define-method new-+ args  (new-+ (car args) (apply new-+ (cdr args)))))
(set! + new-+)

Figure C.3: Extending + for dealing with complex numbers

...
...

Standard primitives in which complex numbers are involved could also be redefined in the same manner. This ends this brief presentation of the STklos extension.

86  STk Reference Manual

Appendix D

Modules: Examples

This appendix shows some usages of the STk modules. Most of the examples which are exhibited here are derived from the Tung and Dybvig paper [5].

Interactive Redefinition

Consider first the definitions,

    (define-module A
      (export square)
      (define square
        (lambda (x) (+ x x))))
    (define-module B
      (import A)
      (define distance
        (lambda (x y)
          (sqrt (+ (square x) (square y))))))

Obviously, the square function exported from A is incorrect, as we can see in its usage below:

    (with-module B (round (distance 3 4)))
              ⟹ 4.0

The function can be redefined (corrected) by the following expression:

    (with-module A
      (set! square
            (lambda (x) (* x x))))

And now,

      (with-module B (round (distance 3 4)))
              ⟹ 5

which is correct.

Lexical principle

This example reuses the modules A and B of previous section and adds a Compare module that exports the less-than-4? predicates, which states if the distance from a point to the origin is less than 4.

(define-module A
(export square)
(define square (lambda (x) (* x x))))
(define-module B
(import A)
(export distance)
(define distance
  (lambda (x y) (sqrt (+ (square x) (square y))))))
(define-module Compare
(import B)
(define less-than-4? (lambda (x y) (< (distance x y) 4)))
(define square       (lambda (x)   (+ x x))))

Consider now the call,

(with-module compare (less-than-4? 3 4))
        ⟹ #f

The call to distance done from less-than-4? indirectly calls the square procedure of module A rather than the one defined locally in module Compare.

Mutually Referential Modules

This example uses two mutually referential modules taht import and export to each other to implement mutually recursive even? and odd? procedures

(define-module Odd)   ;; Forward declaration
(define-module Even
(import Odd)
(export even?)
(define even? (lambda (x) (if (zero? x) #t (odd? (- x 1))))))
(define-module Odd
(import Even)
(export odd?)
(define odd? (lambda (x) (if (zero? x) #f (even? (- x 1))))))

Hereafter are some usages of theses procedures:

(with-module Odd (odd? 3))
          ⟹ #t
(with-module Odd (odd? 10))
          ⟹ #f
(with-module Even (even? 3))
          ⟹ #f
(with-module Even (even? 10))
          ⟹ #t

90  STk Reference Manual

Appendix E

Changes

Introduction

This appendix lists the main differences^{1} among the various recent versions of STk. Differences with older versions as well as implementation changes are described in the CHANGES file located in the main directory of the STk distribution.

Release 3.99.0

Release date: 04/10/98 Changes can be classified in three categories:

• About Scheme – A module system has been added

– Integration of the Bigloo match-case and match-lambda primitives. Furthermore, the file bigloo.stk provides some compatibility between STk and bigloo modules.

– A simple Foreign Function Interface has been added.

– integrates the R^{5}RS values and call-with-values

– multi-line comments have been added.

– new file primitives: remove-file, rename-file and temporary-file-name.
– new list primitives: append!, last-pair, remq, remv and remove.

– load, try-load and autoload? can nw be called with a module as second parameter. If this second parameter is present, the loading is done in the environment of the given module.

• About Tk – Integration of the Tk8.0 toolkit

– Buttons, Checkbuttons and Radiobuttons can use a :variable and :textvariable in a

given environment. This environment is given with the new :environment option.

• About STklos – The MOP of STklos is now very similar to the CLOS’s MOP. In particular generic function has been added for controlling slot accesses, as well as numerous introspection functions.

1Only the differences which affect the language or new ports are reported here. In particular, internal changes, packages written in Scheme or performance enhancement are not discussed here.

– When a class is redefined, the instances and methods which uses it are redefined accordingly, as in CLOS (i.e. if anew slot is added in a class, all its – direct or indirect – instances will have the new slot added dynamically.

Release 3.1.1

Release date: 09/26/96 This release is a bug correction release. It corrects a lot of bugs. A lot of these bugs prevent to install it on some architectures.

Release 3.1

Release date: 07/24/96

• Version of Tk is now at Tk4.1 level.

• STk has been ported on Windows 95 and Windows NT.

• Ports can have a handler which is executed when port becomes readable or writable (see prim-

  itives when-port-readable and when-port-writable.

• Sockets in server mode allow multiple concurrent connection.

• STklos: Two new methods: object-eqv? and object-equal? which are called when applying eqv? or equal? to instances.

• New primitive:setenv!

Release 3.0

Release date: 01/22/96

• Version of Tk is at Tk4.0p2 level.

• Closures are fully supported by Tk. That means that a callback can be now a Scheme closure with its environment. GC problems with closures and usage of the dirty address-of are definitively gone.

• Strings can contain null charters (printing of strings is more friendly in write mode).

• Signals can be redirected to Scheme closures. The end of a GC is seen as a signal.

• Traces on variables are changed (and re-work now): the associated trace must be a thunk.

• New options for some widgets to be more friendly with Scheme world

• STklos: if a method M is defined and if it is already bound to a procedure, the old procedure is called when no method is applicable.

      (define-method car ((x <integer>)) (- x 1))
      (car 10)
                ⟹ 9
      (car (cons 'a 'b))
                ⟹ a

• Small change in the STklos hierarchy. <widget> is now a subclass of <procedure> and its meta

  class is <procedure-metaclass>.

94  STk Reference Manual

Appendix F

Miscellaneous Informations

1 Introduction

This appendix lists a number of things which cannot go elsewhere in this document. The only link between the items listed her is that they should ease your life when using STk.

2 About STk

2.1 Latest release

STk distribution is available on various sites. The original distribution site is kaolin.unice.fr (134.59.132.7). Files are available through anonymous ftp and are located in the /pub/STk directory. Distribution file names have the form STk-x.y.z.tar.gz, where x and y represent the version the release and sub-release numbers of the package.

2.2 Sharing Code

If you have written code that you want to share with the (small) STk community, you can deposit it in the directory /pub/STk/Incoming of kaolin.unice.fr. Mail me a small note when you deposit a file in this directory so I can put in in its definitive place (/pub/STk/Contrib directory contains the contributed code).

2.3 STk Mailing list

There is a mailing list for STk located on kaolin.unice.fr. The intent of this mailing list is to permit to STk users to share experiences, expose problems, submit ideas and … everything which you find interesting (and which is related to STk). To subscribe to the mailing list, simply send a message with the word subscribe in the Subject: field of you mail. Mail must be sent to the following address: stk-request@kaolin.unice.fr To unsubscribe from the mailing list, send a mail at previous email address with the word unsubscribe in the Subject: field. For more information on the mailing list management send a message with the word help in the Subject: field of your mail. In particular, it is possible to find all the messages which have already been sent on the STk mailing list. Subscription/un-subscription/information requests are processed automatically without human intervention. If you something goes wrong, send a mail to eg@unice.fr. Once you have properly subscribe to the mailing list,

• you can send your messages about STk to stk@kaolin.unice.fr, • you will receive all the messages of the mailing list to the email address you used when you

subscribed to the list.

2.4 STk FAQ

Marc Furrer has set up a FAQ for STk. This FAQ is regularly posted on the STk mailing list. It can also be accessed through http://ltiwww.epfl.ch/ furrer/STk/FAQ.html. ASCII version of the

FAQ is available from http://ltiwww.epfl.ch/ furrer/STk/FAQ.txt.

2.5 Reporting a bug

When you find a bug in STk, please send its description to the following address stk-bugs@kaolin.unice.fr. Don’t forget to indicate the version you use and the architecture the system is compiled on. STk version and architecture can be found by using the version and machine-type Scheme primitives. If possible, try to find a small program which exhibit the bug.

3 STk and Emacs

The Emacs family editors can be customized to ease viewing and editing programs of a particular sort. Hints given below enable a fine “integration” of STk in Emacs.

Automatic scheme-mode setting

Emacs mode can be chosen automatically on the file’s name. To edit file ended by .stk or .stklos in Scheme mode, you have to set the Elisp variable auto-mode-alist to control the correspondence between those suffixes and the scheme mode. The simpler way to set this variable consists to add the following lines in your .emacs startup file.

     ;; Add the '.stk' and '.stklos' suffix in the auto-mode-alist Emacs
     ;; variable. Setting this variable permits to automagically place the
     ;; buffer in scheme-mode.
     (setq-default auto-mode-alist (append auto-mode-alist
                                           ("\\.stk$"    . scheme-mode)
                                           ("\\.stklos$" . scheme-mode)))

Using Emacs and CMU Scheme

CMU Scheme package package permits to run the STk interpreter in an Emacs window. Once the package is loaded, you can send text to the inferior STk interpreter from other buffers containing Scheme source. The CMU Scheme package is distributed with Emacs (both FSF-Emacs and Xemacs) and you should have it if you are running this editor. To use the CMU Scheme package with STk, place the following lines in your .emacs startup file.

     ;; Use cmu-scheme rather than xscheme which is launched by default
     ;; whence running 'run-scheme' (xscheme is wired with CScheme)
     (autoload 'run-scheme "cmuscheme" "Run an inferior Scheme" t)
     (setq scheme-program-name "stk")
     (setq inferior-scheme-mode-hook '(lambda() (split-window)))

After having entered those lines in your .emacs file, you can simply run the STk interpreter by typing

     M-x run-scheme

Read the CMU Scheme documentation (or use the describe-mode Elisp command) for a complete description of this package.

Using Emacs and the Ilisp package

Ilisp is another scheme package which allows to run the STk interpreter in an Emacs window. This is a rich package with a lot of nice features. Ilisp comes pre-installed with Xemacs; it has to be installed with FSF Emacs (the last version of Ilisp can be ftp’ed anonymously from ftp.cs.cmu.edu

(128.2.206.173) in the /user/ai/lang/lisp/util/emacs/ilisp directory).

To use the Ilisp package with STk, place the following lines in your .emacs startup file.

     (autoload 'run-ilisp              "ilisp" "Select a new inferior LISP." t)
     (autoload 'stk                    "ilisp" "Run stk in ILISP." t)
     (add-hook 'ilisp-load-hook
               '(lambda ()
                  (require 'completer)
                  ;; Define STk dialect characteristics
                  (defdialect stk "STk Scheme"
                    scheme
                    (setq comint-prompt-regexp "^STk> ")
                    (setq ilisp-program "stk -interactive")
                    (setq comint-ptyp t)
                    (setq comint-always-scroll t)
                    (setq ilisp-last-command "*"))))

After having entered those lines in your .emacs file, you can simply run the STk interpreter by typing

     M-x stk

The Ilisp package comes with a rich documentation which describe how to customize the package.

Other packages

Another way to use STk and Emacs consists to use a special purpose STk mode. You can find two

such modes in the /pub/Contrib directory of kaolin.unice.fr.

3.1 Using the SLIB package with STk

Aubrey Jaffer maintains a package called SLIB which is a portable Scheme library which provides compatibility and utility functions for all standard Scheme implementations. To use this package, you have just to type

     (require "slib")

and follow the instructions given in the SLIB library to use a particular package. Note: SLIB uses also the require/provide mechanism to load components of the library. Once SLIB has been loaded, the standard STk require and provide are overloaded such as if their parameter is a string this is the old STk procedure which is called, and if their parameter is a symbol, this is the SLIB one which is called.

4 Getting information about Scheme

4.1 The R^{4}RS document

R^{4}RS is the document which fully describe the Scheme Programming Language, it can be found in the Scheme repository (see ??) in the directory:

     ftp.cs.indiana.edu:/pub/scheme-repository/doc

Aubrey Jaffer has also translated this document in HTML. A version of this document is available at

  file://swiss-ftp.ai.mit.edu/pub/scm/HTML/r4rs toc.html

4.2 The Scheme Repository

The main site where you can find (many) informations about Scheme is located in the University of Indiana. The Scheme repository is maintained by David Eby. The repository currently consists of the following areas:

• Lots of scheme code meant for benchmarking, library/support, research, education, and fun.

• On-line documents: Machine readable standards documents, standards proposals, various Schemerelated tech reports, conference papers, mail archives, etc.

• Most of the publicly distributable Scheme Implementations.

• Material designed primarily for instruction.

• Freely-distributable promotional or demonstration material for Scheme-related products.

• Utilities (e.g., Schemeweb, SLaTeX).

• Extraneous stuff, extensions, etc. You can access the Scheme repository with

• ftp.cs.indiana.edu:/pub/scheme-repository
• http://www.cs.indiana.edu/scheme-repository/SRhome.html

The Scheme Repository is mirrored in Europe:

• ftp.inria.fr:/lang/Scheme
• faui80.informatik.uni-erlangen.de:/pub/scheme/yorku
• ftp.informatik.uni-muenchen.de:/pub/comp/programming/languages/scheme/scheme-repository

4.3 Usenet newsgroup and other addresses

There is a usenet newsgroup about the Scheme Programming language: comp.lang.scheme. Following addresses contains also material about the Scheme language

• http://www.cs.cmu.edu:8001/Web/Groups/AI/html/faqs/lang/scheme/top.html contains the

Scheme FAQ.

• http://www-swiss.ai.mit.edu/scheme-home.html is the Scheme Home page at MIT
• http://www.ai.mit.edu/projects/su/su.html is the Scheme Underground web page

Bibliography

[1] William Clinger and Jonathan Rees (editors). Revised^{4} Report on the Algorithmic Language Scheme. ACM Lisp Pointers, 4(3), 1991.

[2] John K. Ousterhout. An X11 toolkit based on the Tcl Language. In USENIX Winter Conference, pages 105–115, January 1991.

[3] John K. Ousterhout. Tcl: an embeddable command language. In USENIX Winter Conference, pages 183–192, January 1990.

[4] Erick Gallesio. Extending the STk interpreter. Technical report, I3S CNRS / Université de Nice - Sophia Antipolis, 1997.

[5] Sho-Huan Simon Tung and R. Kent Dybvig. Reliable interactive programming with modules. LISP and Symbolic Computation, 9:343–358, 1996.

[6] Guy L. Steele Jr. Common Lisp: the Language, 2nd Edition. Digital Press, 12 Crosby Drive, Bedford, MA 01730, USA, 1990.

[7] POSIX Committee. System Application Program Interface (API) [C Language]. Information technology—Portable Operating System Interface (POSIX). IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1990.

[8] Manuel Serrano. Bigloo User’s Manual, v1.9b, June 1997.

[9] C. Queinnec and J-M. Geffroy. Partial Evaluation Applied to Symbolic Pattern Matching with Intelligent Backtrack. In et al M. Billaud, editor, Workshop in Static Analysis, number 81–82 in Bigre, Bordeaux (France), September 1992.

[10] A. Wright and B.Duba. Pattern Matching for Scheme. Technical report, Department of Computer Science, Rice University, October 1993.

[11] John K. Ousterhout. Tcl and the Tk toolkit. Addison-Wesley, 1994.

[12] Gregor Kickzales. Tiny-clos. Source available on parcftp.xerox.com in directory /pub/mops, December 1992.

[13] Apple Computer. Dylan: an Object Oriented Dynamic Language. Apple, April 1992.

[14] Chris Hanson. The sos reference manual, version 1.5. in-line documentation of the SOS package. Source available on martigny.ai.mit.edu in /archive/cph directory, March 1993.

[15] Jim de Rivières Gregor Kickzales and Daniel G. Bobrow. The Art of Meta Object Protocol. MIT Press, 1991.

Index

!. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 38 *. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 15 *argc*. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 71 *argv*. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 71 *debug* . . . . . . . . . . . . . . . . . . . . . . . . . . . 36; 61, 71 *gc-verbose* . . . . . . . . . . . . . . . . . . . . . . . . . . . . 71 *help-path*. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 72 *image-path* . . . . . . . . . . . . . . . . . . . . . . . . . . . . 72 *load-path*. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 71 *load-suffixes* . . . . . . . . . . . . . . . . . . . . . . . . . 71 *load-verbose* . . . . . . . . . . . . . . . . . . . . . . . . . . 71 *print-banner* . . . . . . . . . . . . . . . . . . . . . . . . . . 72 *program-name* . . . . . . . . . . . . . . . . . . . . . . . . . . 72 *root* . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 68; 72 *start-withdrawn*. . . . . . . . . . . . . . . . . . . . . . . 72 *stk-library* . . . . . . . . . . . . . . . . . . . . . . . . . . . 72 *tk-patch-level*. . . . . . . . . . . . . . . . . . . . . . . . 72 *tk-version* . . . . . . . . . . . . . . . . . . . . . . . . . . . . 72 +. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 15 -. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 15 /. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 15 :accessor . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 78 :allocation. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 78 :class. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 78 :getter . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 78 :init-keyword . . . . . . . . . . . . . . . . . . . . . . . . . . . 78 :initform . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 78 :instance . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 78 :setter . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 78 :slot-ref . . . . . . . . . . . . . . . . . . . . . . . . . . . . 78; 79 :slot-set! . . . . . . . . . . . . . . . . . . . . . . . . . . . 78; 79 :virtual . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 78 <. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 15 <= . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 15 <object> . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 76 <top> . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 76; 83 =. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 15 >. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 15 >= . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 15

A

abs . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 15 accessor . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 78 acos . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 16 add-signal-handler! . . . . . . . . . . . . . . . . . . . . 40 address-of . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 39 address? . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 39 and . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 9 angle . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 16; 74 append. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 12 append! . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 12 apply . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 20 apropos . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 61 asin . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 16 assoc . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 12 assq . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 12 assv . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 12 atan . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 16 autoload . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 26 autoload? . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 26

B

basename . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 37 begin . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 10 bind . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 69 bindtags . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 69 boolean? . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 11 break . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 39; 69 button. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 29 bye . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 64

C

c. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 40 c-string->string. . . . . . . . . . . . . . . . . . . . . . . . 57 caar . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 12 cadr . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 12

                                             call-with-current-continuation . . . . . . . 20

call-with-input-file . . . . . . . . . . . . . . . . . . . 21

                                             call-with-input-string. . . . . . . . . . . . . . . . . 21
                                             call-with-output-file . . . . . . . . . . . . . . . . . . 21
                                             call-with-output-string . . . . . . . . . . . . . . . 22

call-with-values. . . . . . . . . . . . . . . . . . . . . . . . 91 call/cc . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 20; 21 canonical-path . . . . . . . . . . . . . . . . . . . . . . . . . . 37 car . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 12 case . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 9 catch . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 21 cdddar. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 12 cddddr. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 12 cdr . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 12 ceiling . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 16 char->integer . . . . . . . . . . . . . . . . . . . . . . . . . . . 17 char-alphabetic?. . . . . . . . . . . . . . . . . . . . . . . . 17 char-ci<=? . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 17 char-ci<? . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 17 char-ci=? . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 17 char-ci>=? . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 17 char-ci>? . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 17 char-downcase . . . . . . . . . . . . . . . . . . . . . . . . . . . 18 char-lower-case?. . . . . . . . . . . . . . . . . . . . . . . . 17 char-numeric? . . . . . . . . . . . . . . . . . . . . . . . . . . . 17 char-ready?. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 24 char-upcase. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 17 char-upper-case?. . . . . . . . . . . . . . . . . . . . . . . . 17 char-whitespace?. . . . . . . . . . . . . . . . . . . . . . . . 17 char<=? . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 17 char<?. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 17 char=?. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 17 char>=? . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 17 char>?. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 17 char? . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 17 characters . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 17 chdir . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 38 class . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 75

class-precedence-list . . . . . . . . . . . . . . . . . . 82

class-slots. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 76 close-input-port. . . . . . . . . . . . . . . . . . . . . . . . 24 close-output-port. . . . . . . . . . . . . . . . . . . . . . . 24 close-port . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 27 closure? . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 20 cmu scheme . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 96 complex? . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 15 cond . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 9 cons . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 12 continuation . . . . . . . . . . . . . . . . . . . . . . . . . . . 20; 59 continuation? . . . . . . . . . . . . . . . . . . . . . . . . . . . 20 copy-tree . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 13 cos . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 16 current-error-port . . . . . . . . . . . . . . . . . . . . . 22 current-input-port . . . . . . . . . . . . . . 22; 23, 24 current-module . . . . . . . . . . . . . . . . . . . . . . . . . . 32 current-output-port . . . . . . . . . . . . . 22; 23, 25

D

decompose-file-name . . . . . . . . . . . . . . . . . . . . 37 default slot value . . . . . . . . . . . . . . . . . . . . . 78 define-class . . . . . . . . . . . . . . . . . . . . . . . . . 75; 76 define-external . . . . . . . . . . . . . . . . . . . . . 55; 54 define-generic . . . . . . . . . . . . . . . . . . . . . . . . . . 82 define-macro . . . . . . . . . . . . . . . . . . . . . . . . . . . . 36 define-method . . . . . . . . . . . . . . . . . . . . . . . . . . . 82 define-module . . . . . . . . . . . . . . . . . . . . 30; 32, 33 delay . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 10; 20 denominator . . . . . . . . . . . . . . . . . . . . . . . . . . 16; 73 describe . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 77 detail. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 63 dirname . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 37 display . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 24 do . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 10 dotimes . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 10 dump . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 59 dynamic-wind . . . . . . . . . . . . . . . . . . . . . . . . . . . . 20

E

emacs editor . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 96 environment->list. . . . . . . . . . . . . . . . . . . . . . . 34 environment? . . . . . . . . . . . . . . . . . . . . . . . . . . . . 34 eq . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 13 eq? . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 11; 41, 42 equal . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 13 equal?. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 11 eqv . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 13 eqv? . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 11 error . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 59 eval . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 58 eval-string. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 58 even? . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 15 exact->inexact . . . . . . . . . . . . . . . . . . . . . . . . . . 16 exact?. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 15 exec . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 39 exit . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 63 exp . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 16 expand-file-name. . . . . . . . . . . . . . . . . . . . . . . . 36 expand-heap. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 60 export. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 31 export-all-symbols . . . . . . . . . . . . . . . . . . . . . 32 export-symbol . . . . . . . . . . . . . . . . . . . . . . . . . . . 32 expt . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 16 external-exists?. . . . . . . . . . . . . . . . . . . . . . . . 57

F

faq . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 96 file-exists? . . . . . . . . . . . . . . . . . . . . . . . . . . . . 37 file-is-directory? . . . . . . . . . . . . . . . . . . . . . 37 file-is-executable? . . . . . . . . . . . . . . . . . . . . 37 file-is-readable?. . . . . . . . . . . . . . . . . . . . . . . 37 file-is-regular?. . . . . . . . . . . . . . . . . . . . . . . . 37 file-is-writable?. . . . . . . . . . . . . . . . . . . . . . . 37 find-module. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 31 floor . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 16 fluid-let. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 9; 10 flush . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 25 for-each . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 20 force . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 20 format . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 24; 59

G

garbage collector . . . . . . . . . . . . . . . . . . . . . . . . . . 56 gc . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 59 gc-stats . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 60 gcd . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 16 gensym. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 14 get-internal-info. . . . . . . . . . . . . . . . . . . . . . . 60 get-keyword. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 28 get-output-string . . . . . . . . . . . . . . . . . . . 23; 25 get-signal-handlers . . . . . . . . . . . . . . . . . . . . 41 get-widget-data . . . . . . . . . . . . . . . . . . . . . . . . . 30 getcwd. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 38 getenv. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 39 getpid. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 38 getter . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 78 glob . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 37; 38 global-environment . . . . . . . . . . . . . . . . . . . . . 34

H

hash-table->list. . . . . . . . . . . . . . . . . . . . . . . . 44 hash-table-for-each . . . . . . . . . . . . . . . . . . . . 43 hash-table-get . . . . . . . . . . . . . . . . . . . . . . . . . . 42 hash-table-hash . . . . . . . . . . . . . . . . . . . . . 42; 41 hash-table-map . . . . . . . . . . . . . . . . . . . . . . . . . . 43 hash-table-put! . . . . . . . . . . . . . . . . . . . . . . . . . 42 hash-table-remove! . . . . . . . . . . . . . . . . . . . . . 43 hash-table-stats. . . . . . . . . . . . . . . . . . . . . . . . 44 hash-table?. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 42 help, getting . . . . . . . . . . . . . . . . . . . . . . . . . . . 72

I

if. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 9 ilisp package . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 97 imag-part . . . . . . . . . . . . . . . . . . . . . . . . . . . . 16; 74 import. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 31 inexact->exact . . . . . . . . . . . . . . . . . . . . . . . . . . 16 inexact? . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 15 initial environment . . . . . . . . . . . . . . . . . . . . . . . . 11 input-port?. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 22 input-string-port? . . . . . . . . . . . . . . . . . . . . . 22 inspect . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 61 instance. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 77 integer->char . . . . . . . . . . . . . . . . . . . . . . . . . . . 17 integer? . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 15 ip number. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 51

K

keyword. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 78 keyword->string . . . . . . . . . . . . . . . . . . . . . . . . . 28 keyword? . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 28

L

label . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 29 lambda . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 9 last-pair . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 13 lcm . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 16 length. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 12 let . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 9 let* . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 9 letrec. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 10 list . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 12; 13 list* . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 13 list->string . . . . . . . . . . . . . . . . . . . . . . . . . . . . 19 list->vector . . . . . . . . . . . . . . . . . . . . . . . . . . . . 20 list-ref . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 12 list-tail . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 12 list? . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 12 load . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 26 log . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 16

M

machine-type . . . . . . . . . . . . . . . . . . . . . . . . . . . . 58 macro . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 35; 36 macro-body . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 36 macro-expand . . . . . . . . . . . . . . . . . . . . . . . . . 35; 36 macro-expand-1 . . . . . . . . . . . . . . . . . . . . . . . . . . 35 macro-expansion . . . . . . . . . . . . . . . . . . . . . . . . . . 35 macro?. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 35 magnitude . . . . . . . . . . . . . . . . . . . . . . . . . . . . 16; 74 make . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 77 make-client-socket . . . . . . . . . . . . . . . . . . . . . 51 make-hash-table . . . . . . . . . . . . . . . . . . . . . . . . . 41 make-keyword . . . . . . . . . . . . . . . . . . . . . . . . . . . . 28 make-polar. . . . . . . . . . . . . . . . . . . . . . . . 16; 73, 80 make-rectangular . . . . . . . . . . . . . . . . 16; 73, 80 make-server-socket. . . . . . . . . . . . . . . . . . 52; 51 make-string. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 18 make-vector. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 19 map . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 20 match-case . . . . . . . . . . . . . . . . . . . . . . . . . . . 47; 91 match-lambda . . . . . . . . . . . . . . . . . . . . . . . . . 47; 91 max . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 15 member. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 12 memq . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 12 memv . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 12 menu . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 29 min . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 15 module-environment . . . . . . . . . . . . . . . . . . . . . 34 module-exports . . . . . . . . . . . . . . . . . . . . . . . . . . 33 module-imports . . . . . . . . . . . . . . . . . . . . . . . . . . 33 module-name. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 33 module-symbols . . . . . . . . . . . . . . . . . . . . . . . . . . 33 module? . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 31 modulo. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 15

N

negative? . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 15 newline . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 24 next-method. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 83 not . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 11 null? . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 12 number->string . . . . . . . . . . . . . . . . . . . . . . . . . . 16 number? . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 14 numerator . . . . . . . . . . . . . . . . . . . . . . . . . . . . 16; 73

O

obj. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 13; 39 odd? . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 15 open-file . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 27 open-input-file . . . . . . . . . . . . . . . . . . . . . . . . . 23 open-input-string. . . . . . . . . . . . . . . . . . . . . . . 23 open-output-file. . . . . . . . . . . . . . . . . . . . . . . . 23 open-output-string . . . . . . . . . . . . . . . . . . . . . 23 or. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 9 output-port? . . . . . . . . . . . . . . . . . . . . . . . . . . . . 22 output-string-port? . . . . . . . . . . . . . . . . . . . . 22

P

pair? . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 12 parent-environment . . . . . . . . . . . . . . . . . . . . . 34 pattern matching. . . . . . . . . . . . . . . . . . . . . . . . . . 46 peek-char . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 24 pid. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 49 port->list . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 27 port->sexp-list . . . . . . . . . . . . . . . . . . . . . . . . . 27 port->string . . . . . . . . . . . . . . . . . . . . . . . . . . . . 27 port->string-list. . . . . . . . . . . . . . . . . . . . . . . 27 positive? . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 15 posix.1 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 39 primitive? . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 20 procedure-body . . . . . . . . . . . . . . . . . . . . . . . . . . 21

procedure-environment . . . . . . . . . . . . . . . . . . 34

procedure? . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 20 process-alive? . . . . . . . . . . . . . . . . . . . . . . . . . . 49 process-continue. . . . . . . . . . . . . . . . . . . . . . . . 50 process-error . . . . . . . . . . . . . . . . . . . . . . . . . . . 50 process-exit-status . . . . . . . . . . . . . . . . . . . . 50 process-input . . . . . . . . . . . . . . . . . . . . . . . . . . . 50 process-kill . . . . . . . . . . . . . . . . . . . . . . . . . . . . 50 process-list . . . . . . . . . . . . . . . . . . . . . . . . . . . . 50 process-output . . . . . . . . . . . . . . . . . . . . . . . . . . 50 process-pid. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 49 process-send-signal . . . . . . . . . . . . . . . . . . . . 50 process-stop . . . . . . . . . . . . . . . . . . . . . . . . . . . . 50 process-wait . . . . . . . . . . . . . . . . . . . . . . . . . . . . 50 process? . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 49 promise? . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 20 provide . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 26; 97 provided? . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 26

Q

quasiquote . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 10 quit . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 63 quote . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 8 quotient . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 15

R

r4rs . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 7; 97 random. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 58 rational? . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 15 rationalize . . . . . . . . . . . . . . . . . . . . . . . . . . 16; 73 read . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 24; 8 read-char . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 24 read-from-string. . . . . . . . . . . . . . . . . . . . . . . . 58 read-line . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 24 real-part . . . . . . . . . . . . . . . . . . . . . . . . . . . . 16; 74 real? . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 15 regexp-replace . . . . . . . . . . . . . . . . . . . . . . . . . . 46 regexp-replace-all . . . . . . . . . . . . . . . . . . . . . 46 regexp? . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 45 regular expression . . . . . . . . . . . . . . . . . . . . . . . . . 44 remainder . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 15 remove. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 13 remove-file. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 38 remq . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 13 remv . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 13 rename-file. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 38 repl-display-prompt . . . . . . . . . . . . . . . . . . . . 72 repl-display-result . . . . . . . . . . . . . . . . . . . . 72 report-error . . . . . . . . . . . . . . . . . . . . . . . . . . . . 72 require . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 26; 97 reverse . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 12 root window . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 68 round . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 16 run-process. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 49

S

scheme repository . . . . . . . . . . . . . . . . . . . . . . . . . 98 select-module . . . . . . . . . . . . . . . . . . . . . . . . . . . 33 set! . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 9; 78, 79 set-car! . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 12 set-cdr! . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 12 set-random-seed!. . . . . . . . . . . . . . . . . . . . . . . . 58 set-signal-handler! . . . . . . . . . . . . . . . . . . . . 40 set-widget-data!. . . . . . . . . . . . . . . . . . . . . . . . 30 setenv! . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 39 setter . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 78 sigabrt . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 39 sigalrm . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 39 sigbus. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 40 sigchld . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 39 sigcld. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 40 sigcont . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 39 sigfpe. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 39 sighup. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 39 sigill. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 39 sigint . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 39; 40 sigio . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 40 sigiot. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 39 sigkill . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 39; 40 siglost . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 40 sigpipe . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 39 sigpoll . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 40 sigprof . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 40 sigsegv . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 39 sigstop . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 39; 40 sigsys. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 40 sigterm . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 39 sigtrap . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 39 sigttin . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 39 sigttou . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 39 sigurg. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 40 sigusr1 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 39 sigwinch . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 40 sigxcpu . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 40 sigxfsz . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 40 sin . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 16 slib package . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 97 slot . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 75 slot-ref . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 77 slot-set! . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 77

socket-accept-connection . . . . . . . . . . . . . . 52

socket-down? . . . . . . . . . . . . . . . . . . . . . . . . . . . . 52 socket-dup . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 53 socket-host-address . . . . . . . . . . . . . . . . . . . . 51 socket-host-name. . . . . . . . . . . . . . . . . . . . . . . . 51 socket-input . . . . . . . . . . . . . . . . . . . . . . . . . . . . 51 socket-local-address . . . . . . . . . . . . . . . . . . . 51 socket-output . . . . . . . . . . . . . . . . . . . . . . . . . . . 51 socket-port-number . . . . . . . . . . . . . . . . . . . . . 51 socket-shutdown . . . . . . . . . . . . . . . . . . . . . . . . . 52 socket? . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 51 sort . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 60; 38 split-string . . . . . . . . . . . . . . . . . . . . . . . . . . . . 19 sqrt . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 16 string . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 18; 19 string->list . . . . . . . . . . . . . . . . . . . . . . . . . . . . 18 string->number . . . . . . . . . . . . . . . . . . . . . . . . . . 16 string->regexp . . . . . . . . . . . . . . . . . . . . . . 44; 45 string->symbol . . . . . . . . . . . . . . . . . . . . . . . . . . 14 string->widget . . . . . . . . . . . . . . . . . . . . . . . . . . 29 string-append . . . . . . . . . . . . . . . . . . . . . . . . . . . 18 string-ci<=? . . . . . . . . . . . . . . . . . . . . . . . . . . . . 18 string-ci<?. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 18 string-ci=?. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 18 string-ci>=? . . . . . . . . . . . . . . . . . . . . . . . . . . . . 18 string-ci>?. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 18 string-copy. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 19 string-fill! . . . . . . . . . . . . . . . . . . . . . . . . . . . . 19 string-find? . . . . . . . . . . . . . . . . . . . . . . . . . . . . 19 string-index . . . . . . . . . . . . . . . . . . . . . . . . . . . . 19 string-length . . . . . . . . . . . . . . . . . . . . . . . . . . . 18 string-lower . . . . . . . . . . . . . . . . . . . . . . . . . . . . 19 string-ref . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 18 string-set!. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 18 string-upper . . . . . . . . . . . . . . . . . . . . . . . . . . . . 19 string<=? . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 18 string<? . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 18 string=? . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 18 string>=? . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 18 string>? . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 18 string? . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 18 substring . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 18 symbol->string . . . . . . . . . . . . . . . . . . . . . . . . . . 14 symbol-bound? . . . . . . . . . . . . . . . . . . . . . . . . . . . 35 symbol? . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 14 system. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 38

T

tan . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 16 temporary-file-name . . . . . . . . . . . . . . . . . . . . 38 the-environment . . . . . . . . . . . . . . . . . . . . . . . . . 34 tilde expansion . . . . . . . . . . . . . . . . . . . . . . . . . . . . 36 time . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 61 tk toolkit . . . . . . . . . . . . . . . . . . . . . . . . . . . 7; 29, 59 tk-command. . . . . . . . . . . . . . . . . . . . . . . . 29; 30, 68 tk-command?. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 29 toolkit . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 7; 29, 59 top level environment . . . . . . . . . . . . . . 11; 34, 78 trace-var . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 59 transcript-off . . . . . . . . . . . . . . . . . . . 26; 27, 73 transcript-on . . . . . . . . . . . . . . . . . . . . 26; 27, 73 truncate . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 16 try-load . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 26

U

uncode. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 60 unless . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 9; 11 until . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 11 untrace-var. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 59

V

values. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 91 vector. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 19 vector->list . . . . . . . . . . . . . . . . . . . . . . . . . . . . 20 vector-copy. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 20 vector-fill! . . . . . . . . . . . . . . . . . . . . . . . . . . . . 20 vector-length . . . . . . . . . . . . . . . . . . . . . . . . . . . 19 vector-ref . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 19 vector-resize . . . . . . . . . . . . . . . . . . . . . . . . . . . 20 vector-set!. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 19 vector? . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 19 version . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 58 view . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 61; 63

W

when . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 9 when-port-readable . . . . . . . . . . . . . . . . . . . . . 25 when-port-writable . . . . . . . . . . . . . . . . . . . . . 26 when-socket-ready. . . . . . . . . . . . . . . . . . . . . . . 53 while . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 11 widget . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 29 widget->string . . . . . . . . . . . . . . . . . . . . . . . . . . 29 widget-name. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 30 widget? . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 29 with-input-from-file . . . . . . . . . . . . . . . . . . . 22

with-input-from-string. . . . . . . . . . . . . . . . . 23

with-module. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 32 with-output-to-file . . . . . . . . . . . . . . . . . . . . 22

with-output-to-string . . . . . . . . . . . . . . . . . . 23

write . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 24 write-char . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 24

X

x window system . . . . . . . . . . . . . . . . . . . . . . . . . . . 7

Z

zero? . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 15